11 11 TEX L^{A}TEX

Linux AX25-HOWTO, Amateur Radio.

Terry Dawson, VK2KTJ, terry@perf.no.itg.telstra.com.au v1.5, 17 October 1997

The Linux Operating System is perhaps the only operating system in the world that can boast native and standard support for the AX.25 packet radio protocol utilised by Amateur Radio Operators worldwide. This document aims to describe how to install and configure this support.

Contents

1 Introduction. 3 1.1 Changes from the previous version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Where to obtain new versions of this document. . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 Other related documentation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 The Packet Radio Protocols and Linux. 5 2.1 How it all fits together. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 The AX.25/NetRom/Rose software components. 6

3.1 Finding the kernel, tools and utility packages. . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.1.1 The kernel source: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.1.2 The network tools: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.1.3 The AX25 utilities: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Installing the AX.25/NetRom/Rose software. 8

4.1 Compiling the kernel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.1.1 A word about Kernel modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.1.2 What’s new in 2.0.*+ModuleXX or 2.1.* Kernels ? . . . . . . . . . . . . . . . . . . . . 10

4.2 The network configuration tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2.1 A patch kit that adds Rose support and fixes some bugs. . . . . . . . . . . . . . . . . 11 4.2.2 Building the standard net-tools release. . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4.3 The AX.25 user and utility programs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

5 A note on callsigns, addresses and things before we start. 12 5.1 What are all those T1, T2, N2 and things ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.2 Run time configurable parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

6 Configuring an AX.25 port. 15

6.1 Creating the AX.25 network device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 6.1.1 Creating a KISS device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 6.1.2 Creating a Baycom device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6.1.3 Configuring the AX.25 channel access parameters. . . . . . . . . . . . . . . . . . . . . 18 6.1.4 Creating a SoundModem device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.1.5 Creating a PI card device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 6.1.6 Creating a PacketTwin device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 6.1.7 Creating a generic SCC device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 6.1.8 Creating a BPQ ethernet device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.1.9 Configuring the BPQ Node to talk to the Linux AX.25 support. . . . . . . . . . . . . 30

6.2 Creating the /etc/ax25/axports file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 6.3 Configuring AX.25 routing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

7 Configuring an AX.25 interface for TCP/IP. 32

8 Configuring a NetRom port. 32 8.1 Configuring /etc/ax25/nrports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 8.2 Configuring /etc/ax25/nrbroadcast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 8.3 Creating the NetRom Network device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.4 Starting the NetRom daemon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.5 Configuring NetRom routing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

9 Configuring a NetRom interface for TCP/IP. 35

10 Configuring a Rose port. 36 10.1 Configuring /etc/ax25/rsports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 10.2 Creating the Rose Network device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 10.3 Configuring Rose Routing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

11 Making AX.25/NetRom/Rose calls. 37

12 Configuring Linux to accept Packet connections. 38 12.1 Creating the /etc/ax25/ax25d.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 12.2 A simple example ax25d.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 12.3 Starting ax25d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

13 Configuring the node software. 43 13.1 Creating the /etc/ax25/node.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 13.2 Creating the /etc/ax25/node.perms file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 13.3 Configuring node to run from ax25d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 13.4 Configuring node to run from inetd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

14 Configuring axspawn. 47 14.1 Creating the /etc/ax25/axspawn.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

15 Configuring the pms 48 15.1 Create the /etc/ax25/pms.motd file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 15.2 Create the /etc/ax25/pms.info file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 15.3 Associate AX.25 callsigns with system users. . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 15.4 Add the PMS to the /etc/ax25/ax25d.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . 49 15.5 Test the PMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

16 Configuring the user_call programs. 49

17 Configuring the Rose Uplink and Downlink commands 50 17.1 Configuring a Rose downlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 17.2 Configuring a Rose uplink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 18 Associating AX.25 callsigns with Linux users. 51

19 The /proc/ file system entries. 52

20 AX.25, NetRom, Rose network programming. 53 20.1 The address families. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 20.2 The header files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 20.3 Callsign mangling and examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

21 Some sample configurations. 54 21.1 Small Ethernet LAN with Linux as a router to Radio LAN . . . . . . . . . . . . . . . . . . . 54 21.2 IPIP encapsulated gateway configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 21.3 AXIP encapsulated gateway configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

21.3.1 AXIP configuration options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 21.3.2 A typical /etc/ax25/ax25ipd.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . 60 21.3.3 Running ax25ipd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 21.3.4 Some notes about the routes and route flags . . . . . . . . . . . . . . . . . . . . . . . . 62

21.4 Linking NOS and Linux using a pipe device . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

22 Where do I find more information about .... ? 64 22.1 Packet Radio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 22.2 Protocol Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 22.3 Hardware Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

23 Discussion relating to Amateur Radio and Linux. 64

24 Acknowledgements. 65

25 Copyright. 65

1 Introduction.

This document was originally an appendix to the HAM-HOWTO, but grew too large to be reasonably managed in that fashion. This document describes how to install and configure the native AX.25, NetRom and Rose support for Linux. A few typical configurations are described that could be used as models to work from. The Linux implementation of the amateur radio protocols is very flexible. To people relatively unfamiliar with the Linux operating system the configuration process may look daunting and complicated. It will take you a little time to come to understand how the whole thing fits together. You will find configuration very difficult if you have not properly prepared yourself by learning about Linux in general. You cannot expect to switch from some other environment to Linux without learning about Linux itself.

1.1 Changes from the previous version

Additions:
        Joerg Reuters Web Page
        "More Information" section
        ax25ipd configuration.
Corrections/Updates:
        Changed pty's to a safer range to prevent possible conflicts
        Updated module and ax25-utils versions.
ToDo:
        Fix up the SCC section, this is probably wrong.
        Expand on the programming section.

1.2 Where to obtain new versions of this document.

The best place to obtain the latest version of this document is from a Linux Documentation Project archive. The Linux Documentation Project runs a Web Server and this document appears there as the AX25-HOWTO

<http://sunsite.unc.edu/LDP/HOWTO/AX25-HOWTO.html>. This document is also available in various for-
mats from the sunsite.unc.edu ftp archive <ftp://sunsite.unc.edu/pub/Linux/docs/howto/>.

You can always contact me, but I pass new versions of the document directly to the LDP HOWTO coordinator, so if it isn’t there then chances are I haven’t finished it.

1.3 Other related documentation.

There is a lot of related documentation. There are many documents that relate to Linux networking in more general ways and I strongly recommend you also read these as they will assist you in your efforts and provide you with stronger insight into other possible configurations.

They are:

The HAM-HOWTO <http://sunsite.unc.edu/LDP/HOWTO/HAM-HOWTO.html>,
the NET-3-HOWTO <http://sunsite.unc.edu/LDP/HOWTO/NET-3-HOWTO.html>,
the Ethernet-HOWTO <http://sunsite.unc.edu/LDP/HOWTO/Ethernet-HOWTO.html>,

and:

the Firewall-HOWTO <http://sunsite.unc.edu/LDP/HOWTO/Firewall-HOWTO.html>

More general Linux information may be found by reference to other Linux HOWTO <http://sunsite.unc.

edu/LDP/HOWTO/> documents.

2 The Packet Radio Protocols and Linux.

The AX.25 protocol offers both connected and connectionless modes of operation, and is used either by itself for point-point links, or to carry other protocols such as TCP/IP and NetRom.

It is similar to X.25 level 2 in structure, with some extensions to make it more useful in the amateur radio environment.

The NetRom protocol is an attempt at a full network protocol and uses AX.25 at its lowest layer as a datalink protocol. It provides a network layer that is an adapted form of AX.25. The NetRom protocol features dynamic routing and node aliases.

The Rose protocol was conceived and first implemented by Tom Moulton W2VY and is an implementation of the X.25 packet layer protocol and is designed to operate with AX.25 as its datalink layer protocol. It too provides a network layer. Rose addresses take the form of 10 digit numbers. The first four digits are called the Data Network Identification Code (DNIC) and are taken from Appendix B of the CCITT X.121 recommendation. More information on the Rose protocol may be ontained from the RATS Web server

<http://www.rats.org/>.

Alan Cox developed some early kernel based AX.25 software support for Linux. Jonathon Naylor <g4klx@g4klx.demon.co.uk> has taken up ongoing development of the code, has added NetRom and Rose support and is now the developer of the AX.25 related kernel code. DAMA support was developed by Joerg, DL1BKE, jreuter@poboxes.com. Baycom and SoundModem support were added by Thomas Sailer, <sailer@ife.ee.ethz.ch>. The AX.25 utility software is now maintained by me.

The Linux code supports KISS based TNC’s (Terminal Node Controllers), the Ottawa PI card, the Gracilis PacketTwin card and other Z8530 SCC based cards with the generic SCC driver and both the Parallel and Serial port Baycom modems. Thomas’s new soundmodem driver supports Soundblaster and soundcards based on the Crystal chipset.

The User programs contain a simple PMS (Personal Message System), a beacon facility, a line mode connect program, ‘listen’ an example of how to capture all AX.25 frames at raw interface level and programs to configure the NetRom protocol. Included also are an AX.25 server style program to handle and despatch incoming AX.25 connections and a NetRom daemon which does most of the hard work for NetRom support.

2.1 How it all fits together.

The Linux AX.25 implementation is a brand new implementation. While in many ways it may looks similar to NOS, or BPQ or other AX.25 implementations, it is none of these and is not identical to any of them. The Linux AX.25 implementation is capable of being configured to behave almost identically to other implementations, but the configuration process is very different.

To assist you in understanding how you need to think when configuring this section describes some of the structural features of the AX.25 implementation and how it fits into the context of the overall Linux structure.

Simplified Protocol Layering Diagram

     -----------------------------------------------
     | AF_AX25 | AF_NETROM |  AF_INET    | AF_ROSE |
     |=========|===========|=============|=========|
     |         |           |             |         |
     |         |           |    TCP/IP   |         |

     |         |           ----------    |         |
     |         |   NetRom           |    | Rose    |
     |         -------------------------------------
     |            AX.25                            |
     -----------------------------------------------

This diagram simply illustrates that NetRom, Rose and TCP/IP all run directly on top of AX.25, but that each of these protocols is treated as a seperate protocol at the programming interface. The ‘AF_’ names are simply the names given to the ‘Address Family’ of each of these protocols when writing programs to use them. The important thing to note here is the implicit dependence on the configuration of your AX.25 devices before you can configure your NetRom, Rose or TCP/IP devices.

Software module diagram of Linux Network Implementation

----------------------------------------------------------------------------
 User    | Programs  |   call        node    ||  Daemons | ax25d  mheardd
         |           |   pms         mheard  ||          | inetd  netromd
----------------------------------------------------------------------------
         | Sockets   | open(), close(), listen(), read(), write(), connect()
         |           |------------------------------------------------------
         |           |    AF_AX25   |  AF_NETROM  |   AF_ROSE   |  AF_INET
         |------------------------------------------------------------------
Kernel   | Protocols |    AX.25     |   NetRom    |     Rose    | IP/TCP/UDP
         |------------------------------------------------------------------
         | Devices   |    ax0,ax1   |  nr0,nr1    | rose0,rose1 | eth0,ppp0
         |------------------------------------------------------------------
         | Drivers   |  Kiss   PI2   PacketTwin   SCC   BPQ     | slip ppp
         |           |      Soundmodem      Baycom              | ethernet
----------------------------------------------------------------------------
Hardware | PI2 Card, PacketTwin Card, SCC card, Serial port, Ethernet Card
----------------------------------------------------------------------------

This diagram is a little more general than the first. This diagram attempts to show the relationship between user applications, the kernel and the hardware. It also shows the relationship between the Socket application programming interface, the actual protocol modules, the kernel networking devices and the device drivers. Anything in this diagram is dependent on anything underneath it, and in general you must configure from the bottom of the diagram upwards. So for example, if you want to run the call program you must also configure the Hardware, then ensure that the kernel has the appropriate device driver, that you create the appropriate network device, that the kernel includes the desired protocol that presents a programming interface that the call program can use. I have attempted to lay out this document in roughly that order.

3 The AX.25/NetRom/Rose software components.

The AX.25 software is comprised of three components, the kernel source, the network configuration tools and the utility programs.

The version 2.0.xx Linux kernels include the AX.25, NetRom, Z8530 SCC, PI card and PacketTwin drivers by default. These have been significantly enhanced in the 2.1.* kernels. Unfortunately, the rest of the 2.1.* kernels makes them fairly unstable at the moment and not a good choice for a production system. To solve this problem Jonathon Naylor has prepared a patch kit which will bring the amateur radio protocol support in a 2.0.28 kernel up to the standard of the 2.1.* kernels. This is very simple to apply, and provides a range of facilities not present in the standard kernel such as Rose support.

3.1 Finding the kernel, tools and utility packages.

3.1.1 The kernel source:

The kernel source can be found in its usual place at: ftp.kernel.org

     /pub/linux/kernel/v2.0/linux-2.0.31.tar.gz

The current version of the AX25 upgrade patch is available at: ftp.pspt.fi

     /pub/linux/ham/ax25/ax25-module-14e.tar.gz

3.1.2 The network tools:

The latest alpha release of the standard Linux network tools support AX.25 and NetRom and can be found at: ftp.inka.de

     /pub/comp/Linux/networking/net-tools/net-tools-1.33.tar.gz

The latest ipfwadm package can be found at: ftp.xos.nl

     /pub/linux/ipfwadm/

3.1.3 The AX25 utilities:

There are two different families of AX25-utilities. One is for the 2.0.* kernels and the other will work with either the 2.1.* kernels or the 2.0.*+moduleXX kernels. The ax25-utils version number indicates the oldest version of kernel that they will work with. Please choose a version of the ax25-utils appropriate to your kernel. The following are working combinations. You must use one of the following combinations, any other combination will not work, or will not work well.

     Linux Kernel             AX25 Utility set
     ----------------------   -------------------------
     linux-2.0.29             ax25-utils-2.0.12c.tar.gz **
     linux-2.0.28+module12    ax25-utils-2.1.22b.tar.gz **
     linux-2.0.30+module14c   ax25-utils-2.1.42a.tar.gz
     linux-2.0.31+module14d   ax25-utils-2.1.42a.tar.gz
     linux-2.1.22 ++          ax25-utils-2.1.22b.tar.gz
     linux-2.1.42 ++          ax25-utils-2.1.42a.tar.gz

Note: the ax25-utils-2.0.* series (marked above with the ’**’ symbol) is now obsolete and is no longer supported. This document covers configuration using the versions of software recommended above the table.

While there are differences between the releases, most of the information will be relevant to earlier releases of code.

The AX.25 utility programs can be found at: ftp.pspt.fi <ftp://ftp.pspt.fi/pub/linux/ham/ax25/>

or at: sunsite.unc.edu <ftp://sunsite.unc.edu/pub/Linux/apps/ham/>

4 Installing the AX.25/NetRom/Rose software.

To successfully install AX.25 support on your linux system you must configure and install an appropriate kernel and then install the AX.25 utilities.

4.1 Compiling the kernel.

If you are already familiar with the process of compiling the Linux Kernel then you can skip this section, just be sure to select the appropriate options when compiling the kernel. If you are not, then read on.

The normal place for the kernel source to be unpacked to is the /usr/src directory into a subdirectory called linux. To do this you should be logged in as root and execute a series of commands similar to the following:

     # mv linux linux.old
     # cd /usr/src
     # tar xvfz linux-2.0.31.tar.gz
     # tar xvfz /pub/net/ax25/ax25-module-14e.tar.gz
     # patch -p0 </usr/src/ax25-module-14/ax25-2.0.31-2.1.47-2.diff
     # cd linux

After you have unpacked the kernel source and applied the upgrade, you need to run the configuration script and choose the options that suit your hardware configuration and the options that you wish built into your kernel. You do this by using the command:

     # make menuconfig

You might also try:

     # make config

I’m going to describe the full screen method (menuconfig) because it is easier to move around, but you use whichever you are most comfortable with.

In either case you will be offered a range of options at which you must answer ‘Y’ or ‘N’. (Note you may also answer ‘M’ if you are using modules. For the sake of simplicity I will assume you are not, please make appropriate modifications if you are).

The options most relevant to an AX.25 configuration are:

Code maturity level options  --->
    ...

[*] Prompt for development and/or incomplete code/drivers
...
General setup  --->
...
[*] Networking support
...
Networking options  --->
...
[*] TCP/IP networking
[?] IP: forwarding/gatewaying
...
[?] IP: tunneling
...
[?] IP: Allow large windows (not recommended if <16Mb of memory)
...
[*] Amateur Radio AX.25 Level 2
[?] Amateur Radio NET/ROM
[?] Amateur Radio X.25 PLP (Rose)
...
Network device support  --->
[*] Network device support
...
[*] Radio network interfaces
[?] BAYCOM ser12 and par96 driver for AX.25
[?] Soundcard modem driver for AX.25
[?] Soundmodem support for Soundblaster and compatible cards
[?] Soundmodem support for WSS and Crystal cards
[?] Soundmodem support for 1200 baud AFSK modulation
[?] Soundmodem support for 4800 baud HAPN-1 modulation
[?] Soundmodem support for 9600 baud FSK G3RUH modulation
[?] Serial port KISS driver for AX.25
[?] BPQ Ethernet driver for AX.25
[?] Gracilis PackeTwin support for AX.25
[?] Ottawa PI and PI/2 support for AX.25
[?] Z8530 SCC KISS emulation driver for AX.25
...

The options I have flagged with a ‘*’ are those that you must must answer ‘Y’ to. The rest are dependent on what hardware you have and what other options you want to include. Some of these options are described in more detail later on, so if you don’t know what you want yet, then read ahead and come back to this step later.

After you have completed the kernel configuration you should be able to cleanly compile your new kernel:

 # make dep
 # make clean
 # make zImage

maake sure you move your arch/i386/boot/zImage file wherever you want it and then edit your /etc/lilo.conf file and rerun lilo to ensure that you actually boot from it.

4.1.1 A word about Kernel modules

I recommend that you don’t compile any of the drivers as modules. In nearly all installations you gain nothing but additional complexity. Many people have problems trying to get the modularised components working, not because the software is faulty but because modules are more complicated to install and configure.

If you’ve chosen to compile any of the components as modules, then you’ll also need to use:

     # make modules
     # make modules_install

to install your modules in the appropriate location.

You will also need to add some entries into your /etc/conf.modules file that will ensure that the kerneld program knows how to handle the kernel modules. You should add/modify the following:

     alias net-pf-3     ax25
     alias net-pf-6     netrom
     alias net-pf-11    rose
     alias tty-ldisc-1  slip
     alias tty-ldisc-3  ppp
     alias tty-ldisc-5  mkiss
     alias bc0          baycom
     alias nr0          netrom
     alias pi0a         pi2
     alias pt0a         pt
     alias scc0         optoscc    (or one of the other scc drivers)
     alias sm0          soundmodem
     alias tunl0        newtunnel
     alias char-major-4 serial
     alias char-major-5 serial
     alias char-major-6 lp

4.1.2 What’s new in 2.0.*+ModuleXX or 2.1.* Kernels ?

The 2.1.* kernels have enhanced versions of nearly all of the protocols and drivers. The most significant of the enhancements are:

modularised

the protocols and drivers have all been modularised so that you can insmod and rmmod them whenever you wish. This reduces the kernel memory requirements for infrequently used modules and makes development and bug hunting much simpler. That being said, it also makes configuration slightly more difficult.

All drivers are now network drivers all of the network devices such as Baycom, SCC, PI, Packettwin etc now present a normal network interface, that is they now look like the ethernet driver does, they no longer look like KISS TNC’s. A new utility called net2kiss allows you to build a kiss interface to these devices if you wish.

bug fixed there have been many bug fixes and new features added to the drivers and protocols. The Rose protocol is one important addition.

4.2 The network configuration tools.

Now that you have compiled the kernel you should compile the new network configuration tools. These tools allow you to modify the configuration of network devices and to add routes to the routing table.

The new alpha release of the standard net-tools package includes support for AX.25 and NetRom support. I’ve tested this and it seems to work well for me.

4.2.1 A patch kit that adds Rose support and fixes some bugs.

The standard net-tools-1.33.tar.gz package has some small bugs that affect the AX.25 and NetRom support. I’ve made a small patch kit that corrects these and adds Rose support to the tools as well.

You can get the patch from: zone.pspt.fi <ftp://zone.pspt.fi/pub/linux/ham/ax25/net-tools-1.33.
rose.tjd.diff.gz>.

4.2.2 Building the standard net-tools release.

Don’t forget to read the Release file and follow any instructions there. The steps I used to compile the tools were:

     # cd /usr/src
     # tar xvfz net-tools-1.33.tar.gz
     # zcat net-tools-1.33.rose.tjd.diff.gz | patch -p0
     # cd net-tools-1.33
     # make config

At this stage you will be presented with a series of configuration questions, similar to the kernel configuration questions. Be sure to include support for all of the protocols and network devices types that you intend to use. If you do not know how to answer a particular question then answer ‘Y’.

When the compilation is complete, you should use the:

     # make install

command to install the programs in their proper place.

If you wish to use the IP firewall facilities then you will need the latest firewall administration tool ipfwadm. This tool replaces the older ipfw tool which will not work with new kernels.

I compiled the ipfwadm utility with the following commands:

     # cd /usr/src
     # tar xvfz ipfwadm-2.0beta2.tar.gz
     # cd ipfwadm-2.0beta2
     # make install
     # cp ipfwadm.8 /usr/man/man8
     # cp ipfw.4 /usr/man/man4

4.3 The AX.25 user and utility programs.

After you have successfully compiled and booted your new kernel, you need to compile the user programs. To compile and install the user programs you should use a series of commands similar to the following:

     # cd /usr/src
     # tax xvfz ax25-utils-2.1.42a.tar.gz
     # cd ax25-utils-2.1.42a
     # make config
     # make
     # make install

The files will be installed under the /usr directory by default in subdirectories: bin, sbin, etc and man.

If this is a first time installation, that is you’ve never installed any ax25 utilities on your machine before you should also use the:

     # make installconf

command to install some sample configuration files into the /etc/ax25/ directory from which to work.

If you get messages something like:

gcc -Wall -Wstrict-prototypes -O2 -I../lib -c call.c
call.c: In function `statline':
call.c:268: warning: implicit declaration of function `attron'
call.c:268: `A_REVERSE' undeclared (first use this function)
call.c:268: (Each undeclared identifier is reported only once
call.c:268: for each function it appears in.)

then you should double check that you have the ncurses package properly installed on your system. The configuration script attempts to locate your ncurses packages in the common locations, but some installations have ncurses badly installed and it is unable to locate them.

5 A note on callsigns, addresses and things before we start.

Each AX.25 and NetRom port on your system must have a callsign/ssid allocated to it. These are configured in the configuration files that will be described in detail later on.

Some AX.25 implementations such as NOS and BPQ will allow you to configure the same callsign/ssid on each AX.25 and NetRom port. For somewhat complicated technical reasons Linux does not allow this. This isn’t as big a problem in practise as it might seem.

This means that there are things you should be aware of and take into consideration when doing your configurations.

1. Each AX.25 and NetRom port must be configured with a unique callsign/ssid.

2. TCP/IP will use the callsign/ssid of the AX.25 port it is being transmitted or received by, ie the one you configured for the AX.25 interface in point 1.

3. NetRom will use the callsign/ssid specified for it in its configuration file, but this callsign is only used when your NetRom is speaking to another NetRom, this is not the callsign/ssid that AX.25 users who wish to use your NetRom ‘node’ will use. More on this later.

4. Rose will, by default, use the callsign/ssid of the AX.25 port, unless the Rose callsign has been specifically set using the ‘rsparms’ command. If you set a callsign/ssid using the ‘rsparms’ command then Rose will use this callsign/ssid on all ports.

5. Other programs, such as the ‘ax25d’ program can listen using any callsign/ssid that they wish and these may be duplicated across different ports.

6. If you are careful with routing you can configure the same IP address on all ports if you wish.

5.1 What are all those T1, T2, N2 and things ?

Not every AX.25 implementation is a TNC2. Linux uses nomenclature that differs in some respects from that you will be used to if your sole experience with packet is a TNC. The following table should help you interpret what each of the configurable items are, so that when you come across them later in this text you’ll understand what they mean.

     -------------------------------------------------------------------
     Linux  | TAPR TNC | Description
     -------------------------------------------------------------------
     T1     | FRACK    | How long to wait before retransmitting an
            |          | unacknowledged frame.
     -------------------------------------------------------------------
     T2     | RESPTIME | The minimum amount of time to wait for another
            |          | frame to be received before transmitting
            |          | an acknowledgement.
     -------------------------------------------------------------------
     T3     | CHECK    | The period of time we wait between sending
            |          | a check that the link is still active.
     -------------------------------------------------------------------
     N2     | RETRY    | How many times to retransmit a frame before
            |          | assuming the connection has failed.
     -------------------------------------------------------------------
     Idle   |          | The period of time a connection can be idle
            |          | before we close it down.
     -------------------------------------------------------------------
     Window | MAXFRAME | The maximum number of unacknowledged
            |          | transmitted frames.
     -------------------------------------------------------------------

5.2 Run time configurable parameters

The 2.1.* and 2.0.* +moduleXX kernels have a new feature that allows you to change many previously unchangable parameters at run time. If you take a careful look at the /proc/sys/net/ directory structure you will see many files with useful names that describe various parameters for the network configuration. The files in the /proc/sys/net/ax25/ directory each represents one configured AX.25 port. The name of the file relates to the name of the port.

The structure of the files in /proc/sys/net/ax25/<portname>/ is as follows:

FileName              Meaning              Values                  Default
ip_default_mode       IP Default Mode      0=DG 1=VC               0
ax25_default_mode     AX.25 Default Mode   0=Normal 1=Extended     0
backoff_type          Backoff              0=Linear 1=Exponential  1
connect_mode          Connected Mode       0=No 1=Yes              1
standard_window_size  Standard Window      1  <= N <= 7            2
extended_window_size  Extended Window      1  <= N <= 63           32
t1_timeout            T1 Timeout           1s <= N <= 30s          10s
t2_timeout            T2 Timeout           1s <= N <= 20s          3s
t3_timeout            T3 Timeout           0s <= N <= 3600s        300s
idle_timeout          Idle Timeout         0m <= N                 20m
maximum_retry_count   N2                   1  <= N <= 31           10
maximum_packet_length AX.25 Frame Length   1  <= N <= 512          256

In the table T1, T2 and T3 are given in seconds, and the Idle Timeout is given in minutes. But please note that the values used in the sysctl interface are given in internal units where the time in seconds is multiplied by 10, this allows resolution down to 1/10 of a second. With timers that are allowed to be zero, eg T3 and Idle, a zero value indicates that the timer is disabled.

The structure of the files in /proc/sys/net/netrom/ is as follows:

FileName                       Values                  Default
default_path_quality                                   10
link_fails_count                                       2
network_ttl_initialiser                                16
obsolescence_count_initialiser                         6
routing_control                                        1
transport_acknowledge_delay                            50
transport_busy_delay                                   1800
transport_maximum_tries                                3
transport_requested_window_size                        4
transport_timeout                                      1200

The structure of the files in /proc/sys/net/rose/ is as follows:

FileName                       Values                  Default
acknowledge_hold_back_timeout                          50
call_request_timeout                                   2000
clear_request_timeout                                  1800

link_fail_timeout                                      1200
maximum_virtual_circuits                               50
reset_request_timeout                                  1800
restart_request_timeout                                1800
routing_control                                        1
window_size                                            3

To set a parameter all you need to do is write the desired value to the file itself, for example to check and set the Rose window size you’d use something like:

     # cat /proc/sys/net/rose/window_size
     3
     # echo 4 >/proc/sys/net/rose/window_size
     # cat /proc/sys/net/rose/window_size
     4

6 Configuring an AX.25 port.

Each of the AX.25 applications read a particular configuration file to obtain the parameters for the various AX.25 ports configured on your Linux machine. For AX.25 ports the file that is read is the /etc/ax25/axport file. You must have an entry in this file for each AX.25 port you want on your system.

6.1 Creating the AX.25 network device.

The network device is what is listed when you use the ‘ifconfig’ command. This is the object that the Linux kernel sends and receives network data from. Nearly always the network device has a physical port associated with it, but there are occasions where this isn’t necessary. The network device does relate directly to a device driver.

In the Linux AX.25 code there are a number of device drivers. The most common is probably the KISS driver, but others are the SCC driver(s), the Baycom driver and the SoundModem driver.

Each of these device drivers will create a network device when it is started.

6.1.1 Creating a KISS device.

Kernel Compile Options:

     General setup  --->
         [*] Networking support
     Network device support  --->
         [*] Network device support
         ...
         [*] Radio network interfaces
         [*] Serial port KISS driver for AX.25

Probably the most common configuration will be for a KISS TNC on a serial port. You will need to have the TNC preconfigured and connected to your serial port. You can use a communications program like minicom or seyon to configure the TNC into kiss mode.

To create a KISS device you use the kissattach program. In it simplest form you can use the kissattach program as follows:

     # /usr/sbin/kissattach /dev/ttyS0 radio
     # kissparms -p radio -t 100 -s 100 -r 25

The kissattach command will create a KISS network device. These devices are called ‘ax[0-9]’. The first time you use the kissattach command it creates ‘ax0’, the second time it creates ‘ax1’ etc. Each KISS device has an associated serial port.

The kissparms command allows you to set various KISS parameters on a KISS device.

Specifically the example presented would create a KISS network device using the serial device ‘/dev/ttyS0’ and the entry from the /etc/ax25/axports with a port name of ‘radio’. It then configures it with a txdelay and slottime of 100 milliseconds and a ppersist value of 25.

Please refer to the man pages for more information.

Configuring for Dual Port TNC’s The mkiss utility included in the ax25-utils distribution allows you to make use of both modems on a dual port TNC. Configuration is fairly simple. It works by taking a single serial device connected to a single multiport TNC and making it look like a number of devices each connected to a single port TNC. You do this before you do any of the AX.25 configuration. The devices that you then do the AX.25 configuration on are pseudo-TTY interfaces, (/dev/ttyq*), and not the actual serial device. Pseudo-TTY devices create a kind of pipe through which programs designed to talk to tty devices can talk to other programs designed to talk to tty devices. Each pipe has a master and a slave end. The master end is generally called ‘/dev/ptyq*’ and the slave ends are called ‘/dev/ttyq*’. There is a one to one relationship between masters and slaves, so /dev/ptyq0 is the master end of a pipe with /dev/ttyq0 as its slave. You must open the master end of a pipe before opening the slave end. mkiss exploits this mechanism to split a single serial device into seperate devices.

Example: if you have a dual port tnc and it is connected to your /dev/ttyS0 serial device at 9600 bps, the command:

     # /usr/sbin/mkiss -s 9600 /dev/ttyS0 /dev/ptyq0 /dev/ptyq1
     # /usr/sbin/kissattach /dev/ttyq0 port1
     # /usr/sbin/kissattach /dev/ttyq1 port2

would create two pseudo-tty devices that each look like a normal single port TNC. You would then treat /dev/ttyq0 and /dev/ttyq1 just as you would a conventional serial device with TNC connected. This means you’d then use the kissattach command as described above, on each of those, in the example for AX.25 ports called port1 and port2. You shouldn’t use kissattach on the actual serial device as the mkiss program uses it.

The mkiss command has a number of optional arguments that you may wish to use. They are summarised as follows:

-c enables the addition of a one byte checksum to each KISS frame. This is not supported by most KISS implementation, it is supported by the G8BPG KISS rom.

-s <speed>

sets the speed of the serial port.

-h

enables hardware handshaking on the serial port, it is off by default. Most KISS implementation do not support this, but some do.

-l

enables logging of information to the syslog logfile.

6.1.2 Creating a Baycom device.

Kernel Compile Options:

     Code maturity level options  --->
         [*] Prompt for development and/or incomplete code/drivers
     General setup  --->
         [*] Networking support
     Network device support  --->
         [*] Network device support
         ...
         [*] Radio network interfaces
         [*] BAYCOM ser12 and par96 driver for AX.25

Thomas Sailer, <sailer@ife.ee.ethz.ch>, despite the popularly held belief that it would not work very well, has developed Linux support for Baycom modems. His driver supports the Ser12 serial port, Par96 and the enhanced PicPar parallel port modems. Further information about the modems themselves may be obtained from the Baycom Web site <http://www.baycom.de/>.

Your first step should be to determine the i/o and addresses of the serial or parallel port(s) you have Baycom modem(s) connected to. When you have these you must configure the Baycom driver with them.

The BayCom driver creates network devices called: bc0, bc1, bc2 etc. when it is configured.

The sethdlc utility allows you to configure the driver with these parameters, or, if you have only one Baycom modem installed you may specify the parameters on the insmod commmand line when you load the Baycom module.

For example, a simple configuration. Disable the serial driver for COM1: then configure the Baycom driver for a Ser12 serial port modem on COM1: with the software DCD option enabled:

     # setserial /dev/ttyS0 uart none
     # insmod hdlcdrv
     # insmod baycom mode="ser12*" iobase=0x3f8 irq=4

Par96 parallel port type modem on LPT1: using hardware DCD detection:

     # insmod hdlcdrv
     # insmod baycom mode="par96" iobase=0x378 irq=7 options=0

This is not really the preferred way to do it. The sethdlc utility works just as easily with one device as with many.

The sethdlc man page has the full details, but a couple of examples will illustrate the most important aspects of this configuration. The following examples assume you have already loaded the Baycom module using:

     # insmod hdlcdrv
     # insmod baycom

or that you compiled the kernel with the driver inbuilt.

Configure the bc0 device driver as a Parallel port Baycom modem on LPT1: with software DCD:

     # sethdlc -p -i bc0 mode par96 io 0x378 irq 7

Configure the bc1 device driver as a Serial port Baycom modem on COM1:

     # sethdlc -p -i bc1 mode "ser12*" io 0x3f8 irq 4

6.1.3 Configuring the AX.25 channel access parameters.

The AX.25 channel access parameters are the equivalent of the KISS ppersist, txdelay and slottime type parameters. Again you use the sethdlc utility for this.

Again the sethdlc man page is the source of the most complete information but another example of two won’t hurt:

Configure the bc0 device with TxDelay of 200 mS, SlotTime of 100 mS, PPersist of 40 and half duplex:

     # sethdlc -i bc0 -a txd 200 slot 100 ppersist 40 half

Note that the timing values are in milliseconds.

Configuring the Kernel AX.25 to use the BayCom device The BayCom driver creates standard network devices that the AX.25 Kernel code can use. Configuration is much the same as that for a PI or PacketTwin card.

The first step is to configure the device with an AX.25 callsign. The ifconfig utility may be used to perform this.

     # /sbin/ifconfig bc0 hw ax25 VK2KTJ-15 up

will assign the BayCom device bc0 the AX.25 callsign VK2KTJ-15. Alternatively you can use the axparms command, you’ll still need to use the ifconfig command to bring the device up though:

     # ifconfig bc0 up
     # axparms -setcall bc0 vk2ktj-15

The next step is to create an entry in the /etc/ax25/axports file as you would for any other device. The entry in the axports file is associated with the network device you’ve configured by the callsign you configure.

The entry in the axports file that has the callsign that you configured the BayCom device with is the one that will be used to refer to it.

You may then treat the new AX.25 device as you would any other. You can configure it for TCP/IP, add it to ax25d and run NetRom or Rose over it as you please.

6.1.4 Creating a SoundModem device.

Kernel Compile Options:

     Code maturity level options  --->
         [*] Prompt for development and/or incomplete code/drivers
     General setup  --->
         [*] Networking support
     Network device support  --->
         [*] Network device support
         ...
         [*] Radio network interfaces
         [*] Soundcard modem driver for AX.25
         [?] Soundmodem support for Soundblaster and compatible cards
         [?] Soundmodem support for WSS and Crystal cards
         [?] Soundmodem support for 1200 baud AFSK modulation
         [?] Soundmodem support for 4800 baud HAPN-1 modulation
         [?] Soundmodem support for 9600 baud FSK G3RUH modulation

Thomas Sailer has built a new driver for the kernel that allows you to use your soundcard as a modem. Connect your radio directly to your soundcard to play packet! Thomas recommends at least a 486DX2/66 if you want to use this software as all of the digital signal processing is done by the main CPU.

The driver currently emulates 1200 bps AFSK, 4800 HAPN and 9600 FSK (G3RUH compatible) modem types. The only sound cards currently supported are SoundBlaster and WindowsSoundSystem Compatible models. The sound cards require some circuitry to help them drive the Push-To-Talk circuitry, and information on this is available from Thomas’s SoundModem PTT circuit web page <http:

//www.ife.ee.ethz.ch/~sailer/pcf/ptt_circ/ptt.html>. There are quite a few possible options, they

are: detect the sound output from the soundcard, or use output from a parallel port, serial port or midi port. Circuit examples for each of these are on Thomas’s site.

The SoundModem driver creates network devices called: sm0, sm1, sm2 etc when it is configured.

Note: the SoundModem driver competes for the same resources as the Linux sound driver. So if you wish to use the SoundModem driver you must ensure that the Linux sound driver is not installed. You can of course compile them both as modules and insert and remove them as you wish.

Configuring the sound card. The SoundModem driver does not initialise the sound card. The ax25-utils package includes a utility to do this called ‘setcrystal’ that may be used for SoundCards based on the Crystal chipset. If you have some other card then you will have to use some other software to initialise it. Its syntax is fairly straightforward:

     setcrystal [-w wssio] [-s sbio] [-f synthio] [-i irq] [-d dma] [-c dma2]

So, for example, if you wished to configure a soundblaster card at i/o base address 0x388, irq 10 and DMA 1 you would use:

     # setcrystal -s 0x388 -i 10 -d 1

To configure a WindowSoundSystem card at i/o base address 0x534, irq 5, DMA 3 you would use:

     # setcrystal -w 0x534 -i 5 -d 3

The [-f synthio] parameter is the set the synthesiser address, and the [-c dma2] parameter is to set the second DMA channel to allow full duplex operation.

Configuring the SoundModem driver. When you have configured the soundcard you need to configure the driver telling it where the sound card is located and what sort of modem you wish it to emulate.

The sethdlc utility allows you to configure the driver with these parameters, or, if you have only one soundcard installed you may specify the parameters on the insmod commmand line when you load the SoundModem module.

For example, a simple configuration, with one SoundBlaster soundcard configured as described above emulating a 1200 bps modem:

     # insmod hdlcdrv
     # insmod soundmodem mode="sbc:afsk1200" iobase=0x220 irq=5 dma=1

This is not really the preferred way to do it. The sethdlc utility works just as easily with one device as with many.

The sethdlc man page has the full details, but a couple of examples will illustrate the most important aspects of this configuration. The following examples assume you have already loaded the SoundModem modules using:

     # insmod hdlcdrv
     # insmod soundmodem

or that you compiled the kernel with the driver inbuilt.

Configure the driver to support the WindowsSoundSystem card we configured above to emulate a G3RUH 9600 compatible modem as device sm0 using a parallel port at 0x378 to key the Push-To-Talk:

     # sethdlc -p -i sm0 mode wss:fsk9600 io 0x534 irq 5 dma 3 pario 0x378

Configure the driver to support the SoundBlaster card we configured above to emulate a 4800 bps HAPN modem as device sm1 using the serial port located at 0x2f8 to key the Push-To-Talk:

     # sethdlc -p -i sm1 mode sbc:hapn4800 io 0x388 irq 10 dma 1 serio 0x2f8

Configure the driver to support the SoundBlaster card we configured above to emulate a 1200 bps AFSK modem as device sm1 using the serial port located at 0x2f8 to key the Push-To-Talk:

     # sethdlc -p -i sm1 mode sbc:afsk1200 io 0x388 irq 10 dma 1 serio 0x2f8

Configuring the AX.25 channel access parameters. The AX.25 channel access parameters are the equivalent of the KISS ppersist, txdelay and slottime type parameters. You use the sethdlc utility for this as well.

Again the sethdlc man page is the source of the most complete information but another example of two won’t hurt:

Configure the sm0 device with TxDelay of 100 mS, SlotTime of 50mS, PPersist of 128 and full duplex:

     # sethdlc -i sm0 -a txd 100 slot 50 ppersist 128 full

Note that the timing values are in milliseconds.

Setting the audio levels and tuning the driver. It is very important that the audio levels be set correctly for any radio based modem to work. This is equally true of the SoundModem. Thomas has developed some utility programs that make this task easier. They are called smdiag and smmixer.

smdiag provides two types of display, either an oscilloscope type display or an eye pattern type display.

smmixer allows you to actually adjust the transmit and receive audio levels.

To start the smdiag utility in ’eye’ mode for the SoundModem device sm0 you would use:

     # smdiag -i sm0 -e

To start the smmixer utility for the SoundModem device sm0 you would use:

     # smmixer -i sm0

Configuring the Kernel AX.25 to use the SoundModem The SoundModem driver creates standard network devices that the AX.25 Kernel code can use. Configuration is much the same as that for a PI or PacketTwin card.

The first step is to configure the device with an AX.25 callsign. The ifconfig utility may be used to perform this.

     # /sbin/ifconfig sm0 hw ax25 VK2KTJ-15 up

will assign the SoundModem device sm0 the AX.25 callsign VK2KTJ-15. Alternatively you can use the axparms command, but you still need the ifconfig utility to bring the device up:

     # ifconfig sm0 up
     # axparms -setcall sm0 vk2ktj-15

The next step is to create an entry in the /etc/ax25/axports file as you would for any other device. The entry in the axports file is associated with the network device you’ve configured by the callsign you configure. The entry in the axports file that has the callsign that you configured the SoundModem device with is the one that will be used to refer to it.

You may then treat the new AX.25 device as you would any other. You can configure it for TCP/IP, add it to ax25d and run NetRom or Rose over it as you please.

6.1.5 Creating a PI card device.

Kernel Compile Options:

     General setup  --->
         [*] Networking support
     Network device support  --->
         [*] Network device support
         ...
         [*] Radio network interfaces
         [*] Ottawa PI and PI/2 support for AX.25

The PI card device driver creates devices named ‘pi[0-9][ab]’. The first PI card detected will be allocated ‘pi0’, the second ‘pi1’ etc. The ‘a’ and ‘b’ refer to the first and second physical interface on the PI card. If you have built your kernel to include the PI card driver, and the card has been properly detected then you can use the following command to configure the network device:

     # /sbin/ifconfig pi0a hw ax25 VK2KTJ-15 up

This command would configure the first port on the first PI card detected with the callsign VK2KTJ-15 and make it active. To use the device all you now need to do is to configure an entry into your /etc/ax25/axports file with a matching callsign/ssid and you will be ready to continue on.

The PI card driver was written by David Perry, <dp@hydra.carleton.edu>

6.1.6 Creating a PacketTwin device.

Kernel Compile Options:

     General setup  --->
         [*] Networking support
     Network device support  --->
         [*] Network device support
         ...
         [*] Radio network interfaces
         [*] Gracilis PackeTwin support for AX.25

The PacketTwin card device driver creates devices named ‘pt[0-9][ab]’. The first PacketTwin card detected will be allocated ‘pt0’, the second ‘pt1’ etc. The ‘a’ and ‘b’ refer to the first and second physical interface on the PacketTwin card. If you have built your kernel to include the PacketTwin card driver, and the card has been properly detected then you can use the following command to configure the network device:

     # /sbin/ifconfig pt0a hw ax25 VK2KTJ-15 up

This command would configure the first port on the first PacketTwin card detected with the callsign VK2KTJ-15 and make it active. To use the device all you now need to do is to configure an entry into your /etc/ax25/axports file with a matching callsign/ssid and you will be ready to continue on.

The PacketTwin card driver was written by Craig Small VK2XLZ, <csmall@triode.apana.org.au>.

6.1.7 Creating a generic SCC device.

Kernel Compile Options:

     General setup  --->
         [*] Networking support
     Network device support  --->
         [*] Network device support
         ...
         [*] Radio network interfaces
         [*] Z8530 SCC KISS emulation driver for AX.25

Joerg Reuter, DL1BKE, jreuter@poboxes.com has developed generic support for Z8530 SCC based cards. His driver is configurable to support a range of different types of cards and present an interface that looks like a KISS TNC so you can treat it as though it were a KISS TNC.

Obtaining and building the configuration tool package. While the kernel driver is included in the standard kernel distribution, Joerg distributes more recent versions of his driver with the suite of configuration tools that you will need to obtain as well.

You can obtain the configuration tools package from:

Joerg’s web page <http://www.rat.de/jr/>

or:

db0bm.automation.fh-aachen.de

     /incoming/dl1bke/

or:

insl1.etec.uni-karlsruhe.de

     /pub/hamradio/linux/z8530/

or:

ftp.ucsd.edu

     /hamradio/packet/tcpip/linux
     /hamradio/packet/tcpip/incoming/

You will find multiple versions, choose the one that best suits the kernel you intend to use:

z8530drv-2.4a.dl1bke.tar.gz   2.0.*
z8530drv-utils-3.0.tar.gz    2.1.6 or greater

The following commands were what I used to compile and install the package for kernel version 2.0.30:

     # cd /usr/src
     # gzip -dc z8530drv-2.4a.dl1bke.tar.gz | tar xvpofz -

     # cd z8530drv
     # make clean
     # make dep
     # make module         # If you want to build the driver as a module
     # make for_kernel     # If you want the driver to built into your kernel
     # make install

After the above is complete you should have three new programs installed in your /sbin directory: gencfg, sccinit and sccstat. It is these programs that you will use to configure the driver for your card.

You will also have a group of new special device files created in your /dev called scc0-scc7. These will be used later and will be the ‘KISS’ devices you will end up using.

If you chose to ’make for_kernel’ then you will need to recompile your kernel. To ensure that you include support for the z8530 driver you must be sure to answer ‘Y’ to: ‘Z8530 SCC kiss emulation driver for AX.25’ when asked during a kernel ‘make config’.

If you chose to ’make module’ then the new scc.o will have been installed in the appropriate /lib/modules directory and you do not need to recompile your kernel. Remember to use the insmod command to load the module before your try and configure it.

Configuring the driver for your card. The z8530 SCC driver has been designed to be as flexible as possible so as to support as many different types of cards as possible. With this flexibility has come some cost in configuration.

There is more comprehensive documentation in the package and you should read this if you have any problems. You should particularly look at doc/scc_eng.doc or doc/scc_ger.doc for more detailed information. I’ve paraphrased the important details, but as a result there is a lot of lower level detail that I have not included.

The main configuration file is read by the sccinit program and is called /etc/z8530drv.conf. This file is broken into two main stages: Configuration of the hardware parameters and channel configuration. After you have configured this file you need only add:

     # sccinit

into the rc file that configures your network and the driver will be initialised according to the contents of the configuration file. You must do this before you attempt to use the driver.

Configuration of the hardware parameters. The first section is broken into stanzas, each stanza representing an 8530 chip. Each stanza is a list of keywords with arguments. You may specify up to four SCC chips in this file by default. The #define MAXSCC 4 in scc.c can be increased if you require support for more.

The allowable keywords and arguments are:

chip the chip keyword is used to separate stanzas. It will take anything as an argument. The arguments are not used. data_a

this keyword is used to specify the address of the data port for the z8530 channel ‘A’. The argument is a hexadecimal number e.g. 0x300

ctrl_a this keyword is used to specify the address of the control port for the z8530 channel ‘A’. The arguments is a hexadecimal number e.g. 0x304

data_b this keyword is used to specify the address of the data port for the z8530 channel ‘B’. The argument is a hexadecimal number e.g. 0x301

ctrl_b this keyword is used to specify the address of the control port for the z8530 channel ‘B’. The arguments is a hexadecimal number e.g. 0x305

irq this keyword is used to specify the IRQ used by the 8530 SCC described in this stanza. The argument is an integer e.g. 5

pclock this keyword is used to specify the frequency of the clock at the PCLK pin of the 8530. The argument is an integer frequency in Hz which defaults to 4915200 if the keyword is not supplied.

board the type of board supporting this 8530 SCC. The argument is a character string. The allowed values are:

PA0HZP

the PA0HZP SCC Card EAGLE

the Eagle card PC100

the DRSI PC100 SCC card PRIMUS

the PRIMUS-PC (DG9BL) card BAYCOM

BayCom (U)SCC card

escc this keyword is optional and is used to enable support for the Extended SCC chips (ESCC) such as the 8580, 85180, or the 85280. The argument is a character string with allowed values of ‘yes’ or ‘no’. The default is ‘no’.

vector this keyword is optional and specifies the address of the vector latch (also known as "intack port") for PA0HZP cards. There can be only one vector latch for all chips. The default is 0.

6. Configuring an AX.25 port. 26

special

this keyword is optional and specifies the address of the special function register on several cards. The default is 0.

option

this keyword is optional and defaults to 0.

Some example configurations for the more popular cards are as follows:

BayCom USCC

chip    1
data_a  0x300
ctrl_a  0x304
data_b  0x301
ctrl_b  0x305
irq     5
board   BAYCOM
#
# SCC chip 2
#
chip    2
data_a  0x302
ctrl_a  0x306
data_b  0x303
ctrl_b  0x307
board   BAYCOM

PA0HZP SCC card

chip 1
data_a 0x153
data_b 0x151
ctrl_a 0x152
ctrl_b 0x150
irq 9
pclock 4915200
board PA0HZP
vector 0x168
escc no
#
#
#
chip 2
data_a 0x157
data_b 0x155
ctrl_a 0x156
ctrl_b 0x154
irq 9
pclock 4915200
board PA0HZP

          vector 0x168
          escc no

DRSI SCC card

          chip 1
          data_a 0x303
          data_b 0x301
          ctrl_a 0x302
          ctrl_b 0x300
          irq 7
          pclock 4915200
          board DRSI
          escc no

If you already have a working configuration for your card under NOS, then you can use the gencfg command to convert the PE1CHL NOS driver commands into a form suitable for use in the z8530 driver configuration file.

To use gencfg you simply invoke it with the same parameters as you used for the PE1CHL driver in NET/NOS. For example:

     # gencfg 2 0x150 4 2 0 1 0x168 9 4915200

will generate a skeleton configuration for the OptoSCC card.

Channel Configuration The Channel Configuration section is where you specify all of the other parameters associated with the port you are configuring. Again this section is broken into stanzas. One stanza represents one logical port, and therefore there would be two of these for each one of the hardware parameters stanzas as each 8530 SCC supports two ports.

These keywords and arguments are also written to the /etc/z8530drv.conf file and must appear after the hardware parameters section.

Sequence is very important in this section, but if you stick with the suggested sequence it should work ok. The keywords and arguments are:

device this keyword must be the first line of a port definition and specifies the name of the special device file that the rest of the configuration applies to. e.g. /dev/scc0

speed this keyword specifies the speed in bits per second of the interface. The argument is an integer: e.g.

     1200

clock this keyword specifies where the clock for the data will be sourced. Allowable values are:

dpll normal halfduplex operation external

MODEM supplies its own Rx/Tx clock divider

use fullduplex divider if installed.

mode

this keyword specifies the data coding to be used. Allowable arguments are: nrzi or nrz rxbuffers

this keyword specifies the number of receive buffers to allocate memory for. The argument is an integer, e.g. 8.

txbuffers this keyword specifies the number of transmit buffers to allocate memory for. The argument is an integer, e.g. 8.

bufsize this keyword specifies the size of the receive and transmit buffers. The arguments is in bytes and represents the total length of the frame, so it must also take into account the AX.25 headers and not just the length of the data field. This keyword is optional and default to 384

txdelay

the KISS transmit delay value, the argument is an integer in mS. persist

the KISS persist value, the argument is an integer. slot

the KISS slot time value, the argument is an integer in mS. tail

the KISS transmit tail value, the argument is an integer in mS. fulldup

the KISS full duplex flag, the argument is an integer. 1==Full Duplex, 0==Half Duplex. wait

the KISS wait value, the argument is an integer in mS. min

the KISS min value, the argument is an integer in S. maxkey

the KISS maximum keyup time, the argument is an integer in S. idle

the KISS idle timer value, the argument is an integer in S. maxdef

the KISS maxdef value, the argument is an integer.

group the KISS group value, the argument is an integer.

txoff the KISS txoff value, the argument is an integer in mS.

softdcd the KISS softdcd value, the argument is an integer.

slip the KISS slip flag, the argument is an integer.

Using the driver. To use the driver you simply treat the /dev/scc* devices just as you would a serial tty device with a KISS TNC connected to it. For example, to configure Linux Kernel networking to use your SCC card you could use something like:

     # kissattach -s 4800 /dev/scc0 VK2KTJ

You can also use NOS to attach to it in precisely the same way. From JNOS for example you would use something like:

     attach asy scc0 0 ax25 scc0 256 256 4800

The sccstat and sccparam tools. To assist in the diagnosis of problems you can use the sccstat program to display the current configuration of an SCC device. To use it try:

     # sccstat /dev/scc0

you will displayed a very large amount of information relating to the configuration and health of the

/dev/scc0 SCC port.

The sccparam command allows you to change or modify a configuration after you have booted. Its syntax is very similar to the NOS param command, so to set the txtail setting of a device to 100mS you would use:

     # sccparam /dev/scc0 txtail 0x8

6.1.8 Creating a BPQ ethernet device.

Kernel Compile Options:

     General setup  --->
         [*] Networking support
     Network device support  --->

         [*] Network device support
         ...
         [*] Radio network interfaces
         [*] BPQ Ethernet driver for AX.25

Linux supports BPQ Ethernet compatibility. This enables you to run the AX.25 protocol over your Ethernet LAN and to interwork your linux machine with other BPQ machines on the LAN.

The BPQ network devices are named ‘bpq[0-9]’. The ‘bpq0’ device is associated with the ‘eth0’ device, the ‘bpq1’ device with the ‘eth1’ device etc.

Configuration is quite straightforward. You firstly must have configured a standard Ethernet device. This means you will have compiled your kernel to support your Ethernet card and tested that this works. Refer to the Ethernet-HOWTO <Ethernet-HOWTO.html> for more information on how to do this.

To configure the BPQ support you need to configure the Ethernet device with an AX.25 callsign. The following command will do this for you:

     # /sbin/ifconfig bpq0 hw ax25 vk2ktj-14 up

Again, remember that the callsign you specify should match the entry in the /etc/ax25/axports file that you wish to use for this port.

6.1.9 Configuring the BPQ Node to talk to the Linux AX.25 support.

BPQ Ethernet normally uses a multicast address. The Linux implementation does not, and instead it uses the normal Ethernet broadcast address. The NET.CFG file for the BPQ ODI driver should therefore be modifified to look similar to this:

     LINK SUPPORT
             MAX STACKS 1
             MAX BOARDS 1
     LINK DRIVER E2000                    ; or other MLID to suit your card
             INT 10                       ;
             PORT 300                     ; to suit your card
             FRAME ETHERNET_II
             PROTOCOL BPQ 8FF ETHERNET_II ; required for BPQ - can change PID
     BPQPARAMS                            ; optional - only needed if you want
                                          ; to override the default target addr
             ETH_ADDR  FF:FF:FF:FF:FF:FF  ; Target address

6.2 Creating the /etc/ax25/axports file.

The /etc/ax25/axports is a simple text file that you create with a text editor. The format of the

/etc/ax25/axports file is as follows:
     portname  callsign  baudrate  paclen  window  description

where:

portname is a text name that you will refer to the port by.

callsign is the AX.25 callsign you want to assign to the port.

baudrate is the speed at which you wish the port to communicate with your TNC.

paclen is the maximum packet length you want to configure the port to use for AX.25 connected mode connections.

window is the AX.25 window (K) parameter. This is the same as the MAXFRAME setting of many tnc’s.

description is a textual description of the port.

In my case, mine looks like:

     radio    VK2KTJ-15       4800        256     2       4800bps 144.800 MHz
     ether    VK2KTJ-14       10000000    256     2       BPQ/ethernet device

Remember, you must assign unique callsign/ssid to each AX.25 port you create. Create one entry for each AX.25 device you want to use, this includes KISS, Baycom, SCC, PI, PT and SoundModem ports. Each entry here will describe exactly one AX.25 network device. The entries in this file are associated with the network devices by the callsign/ssid. This is at least one good reason for requiring unique callsign/ssid.

6.3 Configuring AX.25 routing.

You may wish to configure default digipeaters paths for specific hosts. This is useful for both normal AX.25 connections and also IP based connections. The axparms command enables you to do this. Again, the man page offers a complete description, but a simple example might be:

     # /usr/sbin/axparms -route add radio VK2XLZ VK2SUT

This command would set a digipeater entry for VK2XLZ via VK2SUT on the AX.25 port named radio.

7 Configuring an AX.25 interface for TCP/IP.

It is very simple to configure an AX.25 port to carry TCP/IP. If you have KISS interfaces then there are two methods for configuring an IP address. The kissattach command has an option that allows you to do specify an IP address. The more conventional method using the ifconfig command will work on all interface types.

So, modifying the previous KISS example:

     # /usr/sbin/kissattach -i 44.136.8.5 -m 512 /dev/ttyS0 radio
     # /sbin/route add -net 44.136.8.0 netmask 255.255.255.0 ax0
     # /sbin/route add default ax0

to create the AX.25 interface with an IP address of 44.136.8.5 and an MTU of 512 bytes. You should still use the ifconfig to configure the other parameters if necessary.

If you have any other interface type then you use the ifconfig program to configure the ip address and netmask details for the port and add a route via the port, just as you would for any other TCP/IP interface. The following example is for a PI card device, but would work equally well for any other AX.25 network device:

     # /sbin/ifconfig pi0a 44.136.8.5 netmask 255.255.255.0 up
     # /sbin/ifconfig pi0a broadcast 44.136.8.255 mtu 512
     # /sbin/route add -net 44.136.8.0 netmask 255.255.255.0 pi0a
     # /sbin/route add default pi0a

The commands listed above are typical of the sort of configuration many of you would be familiar with if you have used NOS or any of its derivatives or any other TCP/IP software. Note that the default route might not be required in your configuration if you have some other network device configured.

To test it out, try a ping or a telnet to a local host.

     # ping -i 5 44.136.8.58

Note the use of the ‘-i 5’ arguments to ping to tell it to send pings every 5 seconds instead of its default of 1 second.

8 Configuring a NetRom port.

The NetRom protocol relies on, and uses the AX.25 ports you have created. The NetRom protocol rides on top of the AX.25 protocol. To configure NetRom on an AX.25 interface you must configure two files. One file describes the Netrom interfaces, and the other file describes which of the AX.25 ports will carry NetRom. You can configure multiple NetRom ports, each with its own callsign and alias, the same procedure applies for each.

8.1 Configuring /etc/ax25/nrports

The first is the /etc/ax25/nrports file. This file describes the NetRom ports in much the same way as the /etc/ax25/axports file describes the AX.25 ports. Each NetRom device you wish to create must have an entry in the /etc/ax25/nrports file. Normally a Linux machine would have only one NetRom device configured that would use a number of the AX.25 ports defined. In some situations you might wish a special service such as a BBS to have a seperate NetRom alias and so you would create more than one.

This file is formatted as follows:

     name callsign  alias  paclen   description

Where:

name is the text name that you wish to refer to the port by.

callsign is the callsign that the NetRom traffic from this port will use. Note, this is not that address that users should connect to to get access to a node style interface. (The node program is covered later). This callsign/ssid should be unique and should not appear elsewhere in either of the /etc/ax25/axports

     or the /etc/ax25/nrports files.

alias is the NetRom alias this port will have assigned to it.

paclen is the maximum size of NetRom frames transmitted by this port.

description is a free text description of the port.

An example would look something like the following:

     netrom  VK2KTJ-9        LINUX   236     Linux Switch Port

This example creates a NetRom port known to the rest of the NetRom network as ‘LINUX:VK2KTJ-9’.

This file is used by programs such as the call program.

8.2 Configuring /etc/ax25/nrbroadcast

The second file is the /etc/ax25/nrbroadcast file. This file may contain a number of entries. There would normally be one entry for each AX.25 port that you wish to allow NetRom traffic on.

This file is formatted as follows:

     axport min_obs def_qual worst_qual verbose

Where: axport

is the port name obtained from the /etc/ax25/axports file. If you do not have an entry in /etc/ax25/nrbroadcasts for a port then this means that no NetRom routing will occur and any received NetRom broadcasts will be ignored for that port.

min_obs is the minimum obselesence value for the port.

def_qual is the default quality for the port.

worst_qual is the worst quality value for the port, any routes under this quality will be ignored.

verbose is a flag determining whether full NetRom routing broadcasts will occur from this port or only a routing broadcast advertising the node itself.

An example would look something like the following:

     radio    1       200      100         1

8.3 Creating the NetRom Network device

When you have the two configuration files completed you must create the NetRom device in much the same way as you did for the AX.25 devices. This time you use the nrattach command. The nrattach works in just the same way as the axattach command except that it creates NetRom network devices called ‘nr[0-9]’. Again, the first time you use the nrattach command it creates the ‘nr0’ device, the second time it creates the ‘nr1’ network devices etc. To create the network device for the NetRom port we’ve defined we would use:

     # nrattach netrom

This command would start the NetRom device (nr0) named netrom configured with the details specified in

the /etc/ax25/nrports file.

8.4 Starting the NetRom daemon

The Linux kernel does all of the NetRom protocol and switching, but does not manage some functions. The NetRom daemon manages the NetRom routing tables and generates the NetRom routing broadcasts. You start NetRom daemon with the command:

     # /usr/sbin/netromd -i

You should soon see the /proc/net/nr_neigh file filling up with information about your NetRom neighbours.

Remember to put the /usr/sbin/netromd command in your rc files so that it is started automatically each time you reboot.

8.5 Configuring NetRom routing.

You may wish to configure static NetRom routes for specific hosts. The nrparms command enables you to do this. Again, the man page offers a complete description, but a simple example might be:

     # /usr/sbin/nrparms -nodes VK2XLZ-10 + #MINTO 120 5 radio VK2SUT-9

This command would set a NetRom route to #MINTO:VK2XLZ-10 via a neighbour VK2SUT-9 on my AX.25 port called ‘radio’.

You can manually create entries for new neighbours using the nrparms command as well. For example:

     # /usr/sbin/nrparms -routes radio VK2SUT-9 + 120

This command would create VK2SUT-9 as a NetRom neighbour with a quality of 120 and this will be locked and will not be deleted automatically.

9 Configuring a NetRom interface for TCP/IP.

Configuring a NetRom interface for TCP/IP is almost identical to configuring an AX.25 interface for TCP/IP.

Again you can either specify the ip address and mtu on the nrattach command line, or use the ifconfig and route commands, but you need to manually add arp entries for hosts you wish to route to because there is no mechanism available for your machine to learn what NetRom address it should use to reach a particular IP host.

So, to create an nr0 device with an IP address of 44.136.8.5, an mtu of 512 and configured with the details from the /etc/ax25/nrports file for a NetRom port named netrom you would use:

     # /usr/sbin/nrattach -i 44.136.8.5 -m 512 netrom
     # route add 44.136.8.5 nr0

or you could use something like the following commands manually:

     # /usr/sbin/nrattach netrom
     # ifconfig nr0 44.136.8.5 netmask 255.255.255.0 hw netrom VK2KTJ-9
     # route add 44.136.8.5 nr0

Then for each IP host you wish to reach via NetRom you need to set route and arp entries. To reach a destination host with an IP address of 44.136.80.4 at NetRom address BBS:VK3BBS via a NetRom neighbour with callsign VK2SUT-0 you would use commands as follows:

     # route add 44.136.80.4 nr0
     # arp -t netrom -s 44.136.80.4 vk2sut-0
     # nrparms -nodes vk3bbs + BBS 120 6 sl0 vk2sut-0

The ‘120’ and ‘6’ arguments to the nrparms command are the NetRom quality and obsolescence count values for the route.

10 Configuring a Rose port.

The Rose packet layer protocol is similar to layer three of the X.25 specification. The kernel based Rose support is a modified version of the FPAC Rose implementation <http://fpac.lmi.ecp.fr/f1oat/f1oat.

html>.

The Rose packet layer protocol protocol relies on, and uses the AX.25 ports you have created. The Rose protocol rides on top of the AX.25 protocol. To configure Rose you must create a configuration file that describes the Rose ports you want. You can create multiple Rose ports if you wish, the same procedure applies for each.

10.1 Configuring /etc/ax25/rsports

The file where you configure your Rose interfaces is the /etc/ax25/rsports file. This file describes the Rose port in much the same way as the /etc/ax25/axports file describes the AX.25 ports.

This file is formatted as follows:

     name  addresss  description

Where:

name is the text name that you wish to refer to the port by.

address is the 10 digit Rose address you wish to assign to this port.

description is a free text description of the port.

An example would look something like the following:

     rose  5050294760  Rose Port

Note that Rose will use the default callsign/ssid configured on each AX.25 port unless you specify otherwise.

To configure a seperate callsign/ssid for Rose to use on each port you use the rsparms command as follows:

     # /usr/sbin/rsprams -call VK2KTJ-10

This example would make Linux listen for and use the callsign/ssid VK2KTJ-10 on all of the configured AX.25 ports for Rose calls.

10.2 Creating the Rose Network device.

When you have created the /etc/ax25/rsports file you may create the Rose device in much the same way as you did for the AX.25 devices. This time you use the rsattach command. The rsattach command creates network devices named ‘rose[0-5]’. The first time you use the rsattach command it create the ‘rose0’ device, the second time it creates the ‘rose1’ device etc. For example:

     # rsattach rose

This command would start the Rose device (rose0) configured with the details specified in the /etc/ax25/rsports file for the entry named ‘rose’.

10.3 Configuring Rose Routing

The Rose protocol currently supports only static routing. The rsparms utility allows you to configure your Rose routing table under Linux.

For example:

     # rsparms -nodes add 5050295502 radio vk2xlz

would add a route to Rose node 5050295502 via an AX.25 port named ‘radio’ in your /etc/ax25/axports file to a neighbour with the callsign VK2XLZ.

You may specify a route with a mask to capture a number of Rose destinations into a single routing entry. The syntax looks like:

     # rsparms -nodes add 5050295502/4 radio vk2xlz

which would be identical to the previous example except that it would match any destination address that matched the first four digits supplied, in this case any address commencing with the digits 5050. An alternate form for this command is:

     # rsparms -nodes add 5050/4 radio vk2xlz

which is probably the less ambiguous form.

11 Making AX.25/NetRom/Rose calls.

Now that you have all of your AX.25, NetRom and Rose interfaces configured and active, you should be able to make test calls.

The AX25 Utilities package includes a program called ‘call’ which is a splitscreen terminal program for AX.25, NetRom and Rose.

A simple AX.25 call would look like:

     /usr/bin/call radio VK2DAY via VK2SUT

A simple NetRom call to a node with an alias of SUNBBS would look like:

     /usr/bin/call netrom SUNBBS

A simple Rose call to HEARD at node 5050882960 would look like:

     /usr/bin/call rose HEARD 5050882960

Note: you must tell call which port you wish to make the call on, as the same destination node might be reachable on any of the ports you have configured.

The call program is a linemode terminal program for making AX.25 calls. It recognises lines that start with ‘˜ ’ as command lines. The ‘˜ .’ command will close the connection.

Please refer to the man page in /usr/man for more information.

12 Configuring Linux to accept Packet connections.

Linux is a powerful operating system and offers a great deal of flexibility in how it is configured. With this flexibility comes a cost in configuring it to do what you want. When configuring your Linux machine to accept incoming AX.25, NetRom or Rose connections there are a number of questions you need to ask yourself. The most important of which is: "What do I want users to see when they connect?". People are developing neat little applications that may be used to provide services to callers, a simple example is the pms program included in the AX25 utilities, a more complex example is the node program also included in the AX25 utilities. Alternatively you might want to give users a login prompt so that they can make use of a shell account, or you might even have written your own program, such as a customised database or a game, that you want people to connect to. Whatever you choose, you must tell the AX.25 software about this so that it knows what software to run when it accepts an incoming AX.25 connection.

The ax25d program is similar to the inetd program commonly used to accept incoming TCP/IP connections on unix machines. It sits and listens for incoming connections, when it detects one it goes away and checks a configuration file to determine what program to run and connect to that connection. Since this the standard tool for accepting incoming AX.25, NetRom and Rose connections I’ll describe how to configure it.

12.1 Creating the /etc/ax25/ax25d.conf file.

This file is the configuration file for the ax25d AX.25 daemon which handles incoming AX.25, NetRom and Rose connections.

The file is a little cryptic looking at first, but you’ll soon discover it is very simple in practice, with a small trap for you to be wary of.

The general format of the ax25d.conf file is as follows:

     # This is a comment and is ignored by the ax25d program.
     [port_name] || <port_name> || {port_name}
     <peer1>    window T1 T2 T3 idle N2 <mode> <uid> <cmd> <cmd-name> <arguments>
     <peer2>    window T1 T2 T3 idle N2 <mode> <uid> <cmd> <cmd-name> <arguments>
     parameters window T1 T2 T3 idle N2 <mode>

<peer3>    window T1 T2 T3 idle N2 <mode> <uid> <cmd> <cmd-name> <arguments>
   ...
default    window T1 T2 T3 idle N2 <mode> <uid> <cmd> <cmd-name> <arguments>

Where:

# at the start of a line marks a comment and is completely ignored by the ax25d program.

<port_name> is the name of the AX.25, NetRom or Rose port as specified in the /etc/ax25/axports, /etc/ax25/nrports and /etc/ax25/rsports files. The name of the port is surrounded by the ‘[]’ brackets if it is an AX.25 port, the ‘<>’ brackets if it is a NetRom port, or the ‘{}’ brackets if it is a Rose port. There is an alternate form for this field, and that is use prefix the port name with ‘callsign/ssid via’ to indicate that you wish accept calls to the callsign/ssid via this interface. The example should more clearly illustrate this.

<peer> is the callsign of the peer node that this particular configuration applies to. If you don’t specify an SSID here then any SSID will match.

window is the AX.25 Window parameter (K) or MAXFRAME parameter for this configuration.

T1 is the Frame retransmission (T1) timer in half second units.

T2 is the amount of time the AX.25 software will wait for another incoming frame before preparing a response in 1 second units.

T3 is the amount of time of inactivity before the AX.25 software will disconnect the session in 1 second units.

idle is the idle timer value in seconds.

N2 is the number of consecutive retransmissions that will occur before the connection is closed.

<mode> provides a mechanism for determining certain types of general permissions. The modes are enabled or disabled by supplying a combination of characters, each representing a permission. The characters may be in either upper or lower case and must be in a single block with no spaces.

u/U UTMP - currently unsupported. v/V

Validate call - currently unsupported. q/Q

Quiet - Don’t log connection n/N

check NetRom Neighbour - currently unsupported. d/D

Disallow Digipeaters - Connections must be direct, not digipeated. l/L

Lockout - Don’t allow connection. */0

marker - place marker, no mode set.

<uid> is the userid that the program to be run to support the connection should be run as.

<cmd> is the full pathname of the command to be run, with no arguments specified.

<cmd-name> is the text that should appear in a ps as the command name running (normally the same as <cmd> except without the directory path information.

<arguments> are the command line argument to be passed to the <:cmd> when it is run. You pass useful information into these arguments by use of the following tokens:

%d

Name of the port the connection was received on. %U

AX.25 callsign of the connected party without the SSID, in uppercase. %u

AX.25 callsign of the connected party without the SSID, in lowercase. %S

AX.25 callsign of the connected party with the SSID, in uppercase. %s

AX.25 callsign of the connected party with the SSID, in lowercase. %P

AX.25 callsign of the remote node that the connection came in from without the SSID, in uppercase.

%p AX.25 callsign of the remote node that the connection came in from without the SSID, in lowercase.

%R

AX.25 callsign of the remote node that the connection came in from with the SSID, in uppercase. %r

AX.25 callsign of the remote node that the connection came in from with the SSID, in lowercase.

You need one section in the above format for each AX.25, NetRom or Rose interface you want to accept incoming AX.25, NetRom or Rose connections on.

There are two special lines in the paragraph, one starts with the string ‘parameters’ and the other starts with the string ‘default’ (yes there is a difference). These lines serve special functions.

The ‘default’ lines purpose should be obvious, this line acts as a catch-all, so that any incoming connection on the <interface_call> interface that doesn’t have a specific rule will match the ‘default’ rule. If you don’t have a ‘default’ rule, then any connections not matching any specific rule will be disconnected immediately without notice.

The ‘parameters’ line is a little more subtle, and here is the trap I mentioned earlier. In any of the fields for any definition for a peer you can use the ‘*’ character to say ‘use the default value’. The ‘parameters’ line is what sets those default values. The kernel software itself has some defaults which will be used if you don’t specify any using the ‘parameters’ entry. The trap is that the these defaults apply only to those rules below the ‘parameters’ line, not to those above. You may have more than one ‘parameters’ rule per interface definition, and in this way you may create groups of default configurations. It is important to note that the ‘parameters’ rule does not allow you to set the ‘uid’ or ‘command’ fields.

12.2 A simple example ax25d.conf file.

Ok, an illustrative example:

     # ax25d.conf for VK2KTJ - 02/03/97
     # This configuration uses the AX.25 port defined earlier.
     # <peer> Win T1  T2  T3  idl N2 <mode> <uid> <exec> <argv[0]>[<args....>]
     [VK2KTJ-0 via radio]
     parameters 1    10  *  *  *   *   *
     VK2XLZ     *     *  *  *  *   *   *    root  /usr/sbin/axspawn axspawn %u +
     VK2DAY     *     *  *  *  *   *   *    root  /usr/sbin/axspawn axspawn %u +
     NOCALL     *     *  *  *  *   *   L
     default    1    10  5 100 180 5   *    root  /usr/sbin/pms pms -a -o vk2ktj
     [VK2KTJ-1 via radio]
     default    *     *    *   *   *   0    root /usr/sbin/node node
     <netrom>
     parameters 1    10  *  *  *   *   *
     NOCALL     *     *  *  *  *   *   L
     default    *     *  *  *  *   *   0        root /usr/sbin/node node
     {VK2KTJ-0 via rose}
     parameters 1    10  *  *  *   *   *

     VK2XLZ     *     *  *  *  *   *   *    root  /usr/sbin/axspawn axspawn %u +
     VK2DAY     *     *  *  *  *   *   *    root  /usr/sbin/axspawn axspawn %u +
     NOCALL     *     *  *  *  *   *   L
     default    1    10  5 100 180 5   *    root  /usr/sbin/pms pms -a -o vk2ktj
     {VK2KTJ-1 via rose}
     default    *     *    *   *   *   0    root /usr/sbin/node node radio

This example says that anybody attempting to connect to the callsign ‘VK2KTJ-0’ heard on the AX.25 port called ‘radio’ will have the following rules applied:

Anyone whose callsign is set to ‘NOCALL’ should be locked out, note the use of mode ‘L’.

The parameters line changes two parameters from the kernel defaults (Window and T1) and will run the /usr/sbin/axspawn program for them. Any copies of /usr/sbin/axspawn run this way will appear as axspawn in a ps listing for convenience. The next two lines provide definitions for two stations who will receive those permissions.

The last line in the paragraph is the ‘catch all’ definition that everybody else will get (including VK2XLZ and VK2DAY using any other SSID other than -1). This definition sets all of the parameters implicitly and will cause the pms program to be run with a command line argument indicating that it is being run for an AX.25 connection, and that the owner callsign is VK2KTJ. (See the ‘Configuring the PMS’ section below for more details).

The next configuration accepts calls to VK2KTJ-1 via the radio port. It runs the node program for everybody that connects to it.

The next configuration is a NetRom configuration, note the use of the greater-then and less-than braces instead of the square brackets. These denote a NetRom configuration. This configuration is simpler, it simply says that anyone connecting to our NetRom port called ‘netrom’ will have the node program run for them, unless they have a callsign of ‘NOCALL’ in which case they will be locked out.

The last two configurations are for incoming Rose connections. The first for people who have placed calls to ‘vk2ktj-0’ and the second for ‘VK2KTJ-1 at the our Rose node address. These work precisely the same way. Not the use of the curly braces to distinguish the port as a Rose port.

This example is a contrived one but I think it illustrates clearly the important features of the syntax of the configuration file. The configuration file is explained fully in the ax25d.conf man page. A more detailed example is included in the ax25-utils package that might be useful to you too.

12.3 Starting ax25d

When you have the two configuration files completed you start ax25d with the command:

     # /usr/sbin/ax25d

When this is run people should be able to make AX.25 connections to your Linux machine. Remember to put the ax25d command in your rc files so that it is started automatically when you reboot each time.

13. Configuring the node software. 43

13 Configuring the node software.

The node software was developed by Tomi Manninen <tomi.manninen@hut.fi> and was based on the original PMS program. It provides a fairly complete and flexible node capability that is easily configured. It allows users once they are connected to make Telnet, NetRom, Rose, and AX.25 connections out and to obtain various sorts of information such as Finger, Nodes and Heard lists etc. You can configure the node to execute any Linux command you wish fairly simply.

The node would normally be invoked from the ax25d program although it is also capable of being invoked from the TCP/IP inetd program to allow users to telnet to your machine and obtain access to it, or by running it from the command line.

13.1 Creating the /etc/ax25/node.conf file.

The node.conf file is where the main configuration of the node takes place. It is a simple text file and its format is as follows:

# /etc/ax25/node.conf
# configuration file for the node(8) program.
#
# Lines beginning with '#' are comments and are ignored.
# Hostname
# Specifies the hostname of the node machine
hostname        radio.gw.vk2ktj.ampr.org
# Local Network
# allows you to specify what is consider 'local' for the
# purposes of permission checking using nodes.perms.
localnet        44.136.8.96/29
# Hide Ports
# If specified allows you to make ports invisible to users. The
# listed ports will not be listed by the (P)orts command.
hiddenports     rose netrom
# Node Identification.
# this will appear in the node prompt
NodeId          LINUX:VK2KTJ-9
# NetRom port
# This is the name of the netrom port that will be used for
# outgoing NetRom connections from the node.
NrPort          netrom
# Node Idle Timeout
# Specifies the idle time for connections to this node in seconds.
idletimout      1800
# Connection Idle Timeout

     # Specifies the idle timer for connections made via this node in
     # seconds.
     conntimeout     1800
     # Reconnect
     # Specifies whether users should be reconnected to the node
     # when their remote connections disconnect, or whether they
     # should be disconnected complete.
     reconnect       on
     # Command Aliases
     # Provide a way of making complex node commands simple.
     alias           CONV    "telnet vk1xwt.ampr.org 3600"
     alias           BBS     "connect radio vk2xsb"
     # Externam Command Aliases
     # Provide a means of executing external commands under the node.
     # extcmd <cmdname> <flag> <userid> <command>
     # Flag == 1 is the only implemented function.
     # <command> is formatted as per ax25d.conf
     extcmd          PMS     1       root    /usr/sbin/pms pms -u %U -o VK2KTJ
     # Logging
     # Set logging to the system log. 3 is the noisiest, 0 is disabled.
     loglevel        3
     # The escape character
     # 20 = (Control-T)
     EscapeChar      20

13.2 Creating the /etc/ax25/node.perms file.

The node allows you to assign permissions to users. These permissions allow you to determine which users should be allowed to make use of options such as the (T)elnet, and (C)onnect commands, for example, and which shouldn’t. The node.perms file is where this information is stored and contains five key fields. For all fields an asterisk ‘*’ character matches anything. This is useful for building default rules.

user

The first field is the callsign or user to which the permissions should apply. Any SSID value is ignored, so you should just place the base callsign here.

method

Each protocol or access method is also given permissions. For example you might allow users who have connected via AX.25 or NetRom to use the (C)onnect option, but prevent others, such as those who are telnet connected from a non-local node from having access to it. The second field therefore allows you to select which access method this permissions rule should apply to. The access methods allowed are:

          method  description

     ------  -----------------------------------------------------------
     ampr    User is telnet connected from an amprnet address (44.0.0.0)
     ax25    User connected by AX.25
     host    User started node from command line
     inet    user is telnet connected from a non-loca, non-ampr address.
     local   User is telnet connected from a 'local' host
     netrom  User connected by NetRom
     rose    User connected by Rose
     *       User connected by any means.

port

For AX.25 users you can control permissions on a port by port basis too if you choose. This allows you to determine what AX.25 are allowed to do based on which of your ports they have connected to. The third field contains the port name if you are using this facility. This is useful only for AX.25 connections.

password

You may optionally configure the node so that it prompts users to enter a password when they connect. This might be useful to help protect specially configured users who have high authority levels. If the fourth field is set then its value will be the password that will be accepted.

permissions

The permissions field is the final field in each entry in the file. The permissions field is coded as a bit field, with each facility having a bit value which if set allows the option to be used and if not set prevents the facility being used. The list of controllable facilities and their corresponding bit values are:

     value   description
     -----   -------------------------------------------------
      1      Login allowed.
      2      AX25 (C)onnects allowed.
      4      NetRom (C)onnects allowed.
      8      (T)elnet to local hosts allowed.
      16     (T)elnet to amprnet (44.0.0.0) hosts allowed.
      32     (T)elnet to non-local, non-amprnet hosts allowed.
      64     Hidden ports allowed for AX.25 (C)onnects.
      128    Rose (C)onnects allowed.

To code the permissions value for a rule, simply take each of the permissions you want that user to have and add their values together. The resulting number is what you place in field five.

A sample nodes.perms might look like:

# /etc/ax25/node.perms
#
# The node operator is VK2KTJ, has a password of 'secret' and
# is allowed all permissions by all connection methods
vk2ktj  *       *       secret  255
# The following users are banned from connecting

     NOCALL  *       *       *       0
     PK232   *       *       *       0
     PMS     *       *       *       0
     # INET users are banned from connecting.
     *       inet    *       *       0
     # AX.25, NetRom, Local, Host and AMPR users may (C)onnect and (T)elnet
     # to local and ampr hosts but not to other IP addresses.
     *       ax25    *       *       159
     *       netrom  *       *       159
     *       local   *       *       159
     *       host    *       *       159
     *       ampr    *       *       159

13.3 Configuring node to run from ax25d

The node program would normally be run by the ax25d program. To do this you need to add appropriate rules to the /etc/ax25/ax25d.conf file. In my configuration I wanted users to have a choice of either connecting to the node or connecting to other services. ax25d allows you to do this by cleverly creating creating port aliases. For example, given the ax25d configuration presented above, I want to configure node so that all users who connect to VK2KTJ-1 are given the node. To do this I add the following to my

/etc/ax25/ax25d.conf file:
     [vk2ktj-1 via radio]
     default    *     *    *   *   *   0    root /usr/sbin/node node

This says that the Linux kernel code will answer any connection requests for the callsign ‘VK2KTJ-1’ heard on the AX.25 port named ‘radio’, and will cause the node program to be run.

13.4 Configuring node to run from inetd

If you want users to be able to telnet a port on your machine and obtain access to the node you can go this fairly easily. The first thing to decide is what port users should connect to. In this example I’ve arbitrarily chosen port 4000, though Tomi gives details on how you could replace the normal telnet daemon with the node in his documentation.

You need to modify two files.

To /etc/services you should add:

     node    3694/tcp        #OH2BNS's node software

and to /etc/inetd.conf you should add:

     node    stream  tcp     nowait  root    /usr/sbin/node node

When this is done, and you have restarted the inetd program any user who telnet connects to port 3694 of your machine will be prompted to login and if configured, their password and then they will be connected to the node.

14 Configuring axspawn.

The axspawn program is a simple program that allows AX.25 stations who connect to be logged in to your machine. It may be invoked from the ax25d program as described above in a manner similar to the node program. To allow a user to log in to your machine you should add a line similar to the following into your

/etc/ax25/ax25d.conf file:
     default * * * * * 1 root /usr/sbin/axspawn axspawn %u

If the line ends in the + character then the connecting user must hit return before they will be allowed to login. The default is to not wait. Any individual host configurations that follow this line will have the axspawn program run when they connect. When axspawn is run it first checks that the command line argument it is supplied is a legal callsign, strips the SSID, then it checks that /etc/passwd file to see if that user has an account configured. If there is an account, and the password is either "" (null) or + then the user is logged in, if there is anything in the password field the user is prompted to enter a password. If there is not an existing account in the /etc/passwd file then axspawn may be configured to automatically create one.

14.1 Creating the /etc/ax25/axspawn.conf file.

You can alter the behaviour of axspawn in various ways by use of the /etc/ax25/axspawn.conf file. This file is formatted as follows:

     # /etc/ax25/axspawn.conf
     #
     # allow automatic creation of user accounts
     create    yes
     #
     # guest user if above is 'no' or everything else fails. Disable with "no"
     guest     no
     #
     # group id or name for autoaccount
     group     ax25
     #
     # first user id to use
     first_uid 2001
     #
     # maximum user id
     max_uid   3000
     #
     # where to add the home directory for the new users
     home      /home/ax25
     #
     # user shell
     shell     /bin/bash
     #
     # bind user id to callsign for outgoing connects.
     associate yes

The eight configurable characteristics of axspawn are as follows:

# indicates a comment.

create if this field is set to yes then axspawn will attempt to automatically create a user account for any user who connects and does not already have an entry in the /etc/passwd file.

guest this field names the login name of the account that will be used for people who connect who do not already have accounts if create is set to no. This is usually ax25 or guest.

group this field names the group name that will be used for any users who connect and do not already have an entry in the /etc/passwd file.

first_uid this is the number of the first userid that will be automatically created for new users.

max_uid this is the maximum number that will be used for the userid of new users.

home this is the home (login) directory of new users.

shell this is the login shell of any new users.

associate this flag indicates whether outgoing AX.25 connections made by this user after they login will use their own callsign, or your stations callsign.

15 Configuring the pms

The pms program is an implementation of a simple personal message system. It was originally written by Alan Cox. Dave Brown, N2RJT, <dcb@vectorbd.com> has taken on further development of it. At present it is still very simple, supporting only the ability to send mail to the owner of the system and to obtain some limited system information but Dave is working to expand its capability to make it more useful.

After that is done there are a couple of simple files that you should create that give users some information about the system and then you need to add appropriate entries into the ax25d.conf file so that connected users are presented with the PMS.

15.1 Create the /etc/ax25/pms.motd file.

The /etc/ax25/pms.motd file contains the ‘message of the day’ that users will be presented with after they connect and receive the usual BBS id header. The file is a simple text file, any text you include in this file will be sent to users.

15.2 Create the /etc/ax25/pms.info file.

The /etc/ax25/pms.info file is also a simple text file in which you would put more detailed information about your station or configuration. This file is presented to users in response to their issuing of the Info command from the PMS> prompt.

15.3 Associate AX.25 callsigns with system users.

When a connected user sends mail to an AX.25 callsign, the pms expects that callsign to be mapped, or associated with a real system user on your machine. This is described in a section of its own.

15.4 Add the PMS to the /etc/ax25/ax25d.conf file.

Adding the pms to your ax25d.conf file is very simple. There is one small thing you need to think about though. Dave has added command line arguments to the PMS to allow it to handle a number of different text end-of-line conventions. AX.25 and NetRom by convention expect the end-of-line to be carriage return, linefeed while the standard unix end-of-line is just newline. So, for example, if you wanted to add an entry that meant that the default action for a connection received on an AX.25 port is to start the PMS then you would add a line that looked something like:

     default  1  10 5 100 5   0    root  /usr/sbin/pms pms -a -o vk2ktj

This simply runs the pms program, telling it that it is an AX.25 connection it is connected to and that the PMS owner is vk2ktj. Check the man page for what you should specify for other connection methods.

15.5 Test the PMS.

To test the PMS, you can try the following command from the command line:

# /usr/sbin/pms -u vk2ktj -o vk2ktj

Substitute your own callsign for mine and this will run the pms, telling it that it is to use the unix end-of-line convention, and that user logging in is vk2ktj. You can do all the things connected users can.

Additionally you might try getting some other node to connect to you to confirm that your ax25d.conf configuration works.

16 Configuring the user_call programs.

The ‘user_call’ programs are really called: ax25_call and netrom_call. They are very simple programs designed to be called from ax25d to automate network connections to remote hosts. They may of course be called from a number of other places such as shell scripts or other daemons such as the node program.

They are like a very simple call program. They don’t do any meddling with the data at all, so the end of line handling you’ll have to worry about yourself.

Let’s start with an example of how you might use them. Imagine you have a small network at home and that you have one linux machine acting as your Linux radio gateway and another machine, lets say a BPQ node connected to it via an ethernet connection.

Normally if you wanted radio users to be able to connect to the BPQ node they would either have to digipeat through your linux node, or connect to the node program on your linux node and then connect from it. The ax25_call program can simplify this if it is called from the ax25d program.

Imagine the BPQ node has the callsign VK2KTJ-9 and that the linux machine has the AX.25/ethernet port named ‘bpq’. Let us also imagine the Linux gateway machine has a radio port called ‘radio’.

An entry in the /etc/ax25/ax25d.conf that looked like:

     [VK2KTJ-1 via radio]
     default    * * * *   *   *  *
                     root /usr/sbin/ax25_call ax25_call bpq %u vk2ktj-9

would enable users to connect direct to ‘VK2KTJ-1’ which would actually be the Linux ax25d daemon and then be automatically switched to an AX.25 connection to ‘VK2KTJ-9’ via the ‘bpq’ interface.

There are all sorts of other possible configurations that you might try. The ‘netrom_call’ and ‘rose_call’ utilities work in similar ways. One amateur has used this utility to make connections to a remote BBS easier. Normally the users would have to manually enter a long connection string to make the call so he created an entry that made the BBS appear as though it were on the local network by having his ax25d proxy the connection to the remote machine.

17 Configuring the Rose Uplink and Downlink commands

If you are familiar with the ROM based Rose implementation you will be familiar with the method by which AX.25 users make calls across a Rose network. If a users local Rose node has the callsign VK2KTJ-5 and the AX.25 user wants to connect to VK5XXX at remote Rose node 5050882960 then they would issue the command:

     c vk5xxx v vk2ktj-5 5050 882960

At the remote node, VK5XXX would see an incoming connection with the local AX.25 users callsign and being digipeated via the remote Rose nodes callsign.

The Linux Rose implementation does not support this capability in the kernel, but there are two application programs called rsuplnk and rsdwnlnk which perform this function.

17.1 Configuring a Rose downlink

To configure your Linux machine to accept a Rose connection and establish an AX.25 connection to any destination callsign that is not being listened for on your machine you need to add an entry to your /etc/ax25/ax25d.conf file. Normally you would configure this entry to be the default behaviour for incoming Rose connections. For example you might have Rose listeners operating for destinations like NODE-0 or HEARD-0 that you wish to handle locally, but for all other destination calls you may want to pass them to the rsdwnlink command and assume they are AX.25 users.

A typical configuration would look like:

     #
     {* via rose}
     NOCALL   * * * * * *  L
     default  * * * * * *  - root  /usr/sbin/rsdwnlnk rsdwnlnk 4800 vk2ktj-5
     #

With this configuration any user who established a Rose connection to your Linux nodes address with a destination call of something that you were not specifically listening for would be converted into an AX.25 connection on the AX.25 port named 4800 with a digipeater path of VK2KTJ-5.

17.2 Configuring a Rose uplink

To configure your Linux machine to accept AX.25 connections in the same way that a ROM Rose node would you must add an entry into your /etc/ax25/ax25d.conf file that looks similar to the following:

     #
     [VK2KTJ-5* via 4800]
     NOCALL   * * * * * *  L
     default  * * * * * *  - root  /usr/sbin/rsuplnk rsuplnk rose
     #

Note the special syntax for the local callsign. The ‘*’ character indicates that the application should be invoked if the callsign is heard in the digipeater path of a connection.

This configuration would allow an AX.25 user to establish Rose calls using the example connect sequence presented in the introduction. Anybody attempting to digipeat via VK2KTJ-5 on the AX.25 port named 4800 would be handled by the rsuplnk command.

18 Associating AX.25 callsigns with Linux users.

There are a number of situations where it is highly desirable to associate a callsign with a linux user account. One example might be where a number of amateur radio operators share the same linux machine and wish to use their own callsign when making calls. Another is the case of PMS users wanting to talk to a particular user on your machine.

The AX.25 software provides a means of managing this association of linux user account names with callsigns. We’ve mentioned it once already in the PMS section, but I’m spelling it out here to be sure you don’t miss it.

You make the association with the axparms command. An example looks like:

     # axparms -assoc vk2ktj terry

This command associates that AX.25 callsign vk2ktj with the user terry on the machine. So, for example, any mail for vk2ktj on the pms will be sent to Linux account terry.

Remember to put these associations into your rc file so that they are available each time your reboot.

Note you should never associate a callsign with the root account as this can cause configuration problems in other programs.

19 The /proc/ file system entries.

The /proc filesystem contains a number of files specifically related to the AX25 and NetRom kernel software. These files are normally used by the AX52 utilities, but they are plainly formatted so you may be interested in reading them. The format is fairly easily understood so I don’t think much explanation will be necessary.

/proc/net/arp contains the list of Address Resolution Protocol mappings of IP addresses to MAC layer protocol addresses. These can can AX.25, ethernet or some other MAC layer protocol.

/proc/net/ax25

contains a list of AX.25 sockets opened. These might be listening for a connection, or active sessions. /proc/net/ax25_bpqether

contains the AX25 over ethernet BPQ style callsign mappings. /proc/net/ax25_calls

contains the linux userid to callsign mappings set my the axparms -assoc command. /proc/net/ax25_route

contains AX.25 digipeater path information. /proc/net/nr

contains a list of NetRom sockets opened. These might be listening for a connection, or active sessions. /proc/net/nr_neigh

contains information about the NetRom neighbours known to the NetRom software. /proc/net/nr_nodes

contains information about the NetRom nodes known to the NetRom software. /proc/net/rose

contains a list of Rose sockets opened. These might be listening for a connection, or active sessions. /proc/net/rose_nodes

contains a mapping of Rose destinations to Rose neighbours. /proc/net/rose_neigh

contains a list of known Rose neighbours. /proc/net/rose_routes

contains a list of all established Rose connections.

20 AX.25, NetRom, Rose network programming.

Probably the biggest advantage of using the kernel based implementations of the amateur packet radio protocols is the ease with which you can develop applications and programs to use them.

While the subject of Unix Network Programming is outside the scope of this document I will describe the elementary details of how you can make use of the AX.25, NetRom and Rose protocols within your software.

20.1 The address families.

Network programming for AX.25, NetRom and Rose is quite similar to programming for TCP/IP under Linux. The major differences being the address families used, and the address structures that need to be mangled into place.

The address family names for AX.25, NetRom and Rose are AF_AX25, AF_NETROM and AF_ROSE respectively.

20.2 The header files.

You must always include the ‘ax25.h’ header file, and also the ‘netrom.h’ or ‘rose.h’ header files if you are dealing with those protocols. Simple top level skeletons would look something like the following:

For AX.25:

     #include <ax25.h>
     int s, addrlen = sizeof(struct full_sockaddr_ax25);
     struct full_sockaddr_ax25 sockaddr;
     sockaddr.fsa_ax25.sax25_family = AF_AX25

For NetRom:

     #include <ax25.h>
     #include <netrom.h>
     int s, addrlen = sizeof(struct full_sockaddr_ax25);
     struct full_sockaddr_ax25 sockaddr;
     sockaddr.fsa_ax25.sax25_family = AF_NETROM;

For Rose:

     #include <ax25.h>
     #include <rose.h>
     int s, addrlen = sizeof(struct sockaddr_rose);
     struct sockaddr_rose sockaddr;
     sockaddr.srose_family = AF_ROSE;

20.3 Callsign mangling and examples.

There are routines within the lib/ax25.a library built in the AX25 utilities package that manage the callsign conversions for you. You can write your own of course if you wish.

The user_call utilities are excellent examples from which to work. The source code for them is included in the AX25 utilities package. If you spend a little time working with those you will soon see that ninety percent of the work is involved in just getting ready to open the socket. Actually making the connection is easy, the preparation takes time.

The example are simple enough to not be very confusing. If you have any questions, you should feel to direct them to the linux-hams mailing list and someone there will be sure to help you.

21 Some sample configurations.

Following are examples of the most common types of configurations. These are guides only as there are as many ways of configuring your network as there are networks to configure, but they may give you a start.

21.1 Small Ethernet LAN with Linux as a router to Radio LAN

Many of you may have small local area networks at home and want to connect the machines on that network to your local radio LAN. This is the type of configuration I use at home. I arranged to have a suitable block of addresses allocated to me that I could capture in a single route for convenience and I use these on my Ethernet LAN. Your local IP coordinator will assist you in doing this if you want to try it as well. The addresses for the Ethernet LAN form a subset of the radio LAN addresses. The following configuration is the actual one for my linux router on my network at home:

                                               .      .   .    .    . .
       ---                                .
        | Network       /---------\     .    Network
        | 44.136.8.96/29|         |    .     44.136.8/24        \ | /
        |               | Linux   |   .                          \|/
        |               |         |  .                            |
        |          eth0 | Router  |  .  /-----\    /----------\   |
        |---------------|         |-----| TNC |----| Radio    |---/
        |   44.136.8.97 |  and    |  .  \-----/    \----------/
        |               |         | sl0
        |               | Server  | 44.136.8.5
        |               |         |    .
        |               |         |     .
        |               \_________/       .
       ---                                     .      .   .    .    . .
     #!/bin/sh
     # /etc/rc.net
     # This configuration provides one KISS based AX.25 port and one
     # Ethernet device.
     echo "/etc/rc.net"
     echo "  Configuring:"
     echo -n "    loopback:"
     /sbin/ifconfig lo 127.0.0.1

21. Some sample configurations. 55

/sbin/route add 127.0.0.1
echo " done."
echo -n "    ethernet:"
/sbin/ifconfig eth0 44.136.8.97 netmask 255.255.255.248 \
                broadcast 44.136.8.103 up
/sbin/route add 44.136.8.97 eth0
/sbin/route add -net 44.136.8.96 netmask 255.255.255.248 eth0
echo " done."
echo -n "    AX.25: "
kissattach -i 44.136.8.5 -m 512 /dev/ttyS1 4800
ifconfig sl0 netmask 255.255.255.0 broadcast 44.136.8.255
route add -host 44.136.8.5 sl0
route add -net 44.136.8.0 window 1024 sl0
echo -n "    Netrom: "
nrattach -i 44.136.8.5 netrom
echo "  Routing:"
/sbin/route add default gw 44.136.8.68 window 1024 sl0
echo "    default route."
echo done.
# end
/etc/ax25/axports
# name  callsign        speed   paclen  window  description
4800    VK2KTJ-0        4800    256     2       144.800 MHz
/etc/ax25/nrports
# name  callsign        alias   paclen  description
netrom  VK2KTJ-9        LINUX   235     Linux Switch Port
/etc/ax25/nrbroadcast
# ax25_name     min_obs def_qual        worst_qual      verbose
4800            1       120             10              1

• You must have IP_FORWARDING enabled in your kernel.

• The AX.25 configuration files are pretty much those used as examples in the earlier sections, refer to those where necessary.

• I’ve chosen to use an IP address for my radio port that is not within my home network block. I needn’t have done so, I could have easily used 44.136.8.97 for that port too.

• 44.136.8.68 is my local IPIP encapsulated gateway and hence is where I point my default route.

• Each of the machines on my Ethernet network have a route:

     route add -net 44.0.0.0 netmask 255.0.0.0 \
             gw 44.136.8.97 window 512 mss 512 eth0

The use of the mss and window parameters means that I can get optimum performance from both local Ethernet and radio based connections.

• I also run my smail, http, ftp and other daemons on the router machine so that it needs to be the only machine to provide others with facilities.

• The router machine is a lowly 386DX20 with a 20Mb harddrive and a very minimal linux configuration.

21.2 IPIP encapsulated gateway configuration.

Linux is now very commonly used for TCP/IP encapsulated gateways around the world. The new tunnel driver supports multiple encapsulated routes and makes the older ipip daemon obsolete.

A typical configuration would look similar to the following.

                                          .      .   .    .    . .
  ---                                .
   | Network       /---------\     .    Network
   | 154.27.3/24   |         |    .     44.136.16/24       \ | /
   |               | Linux   |   .                          \|/
   |               |         |  .                            |
   |          eth0 | IPIP    |  .  /-----\    /----------\   |
---|---------------|         |-----| TNC |----| Radio    |---/
   |   154.27.3.20 | Gateway |  .  \-----/    \----------/
   |               |         | sl0
   |               |         | 44.136.16.1
   |               |         |    .
   |               |         |     .
   |               \_________/       .
  ---                                     .      .   .    .    . .

The configuration files of interest are:

# /etc/rc.net
# This file is a simple configuration that provides one KISS AX.25
# radio port, one Ethernet device, and utilises the kernel tunnel driver
# to perform the IPIP encapsulation/decapsulation
#
echo "/etc/rc.net"
echo "  Configuring:"
#
echo -n "    loopback:"
/sbin/ifconfig lo 127.0.0.1
/sbin/route add 127.0.0.1
echo " done."
#
echo -n "    ethernet:"
/sbin/ifconfig eth0 154.27.3.20 netmask 255.255.255.0 \

                broadcast 154.27.3.255 up
/sbin/route add 154.27.3.20 eth0
/sbin/route add -net 154.27.3.0 netmask 255.255.255.0 eth0
echo " done."
#
echo -n "    AX.25: "
kissattach -i 44.136.16.1 -m 512 /dev/ttyS1 4800
/sbin/ifconfig sl0 netmask 255.255.255.0 broadcast 44.136.16.255
/sbin/route add -host 44.136.16.1 sl0
/sbin/route add -net 44.136.16.0 netmask 255.255.255.0 window 1024 sl0
#
echo -n "    tunnel:"
/sbin/ifconfig tunl0 44.136.16.1 mtu 512 up
#
echo done.
#
echo -n "Routing ... "
source /etc/ipip.routes
echo done.
#
# end.

and:

# /etc/ipip.routes
# This file is generated using the munge script
#
/sbin/route add -net 44.134.8.0 netmask 255.255.255.0 tunl0 gw 134.43.26.1
/sbin/route add -net 44.34.9.0 netmask 255.255.255.0 tunl0 gw 174.84.6.17
/sbin/route add -net 44.13.28.0 netmask 255.255.255.0 tunl0 gw 212.37.126.3
   ...
   ...
   ...
/etc/ax25/axports
# name  callsign        speed   paclen  window  description
4800    VK2KTJ-0        4800    256     2       144.800 MHz

Some points to note here are:

• The new tunnel driver uses the gw field in the routing table in place of the pointopoint parameter to specify the address of the remote IPIP gateway. This is why it now supports multiple routes per interface.

• You can configure two network devices with the same address. In this example both the sl0 and the tunl0 devices have been configured with the IP address of the radio port. This is done so that the remote gateway sees the correct address from your gateway in encapsulated datagrams sent to it.

• The route commands used to specify the encapsulated routes can be automatically generated by a modified version of the munge script. This is included below. The route commands would then be written to a separate file and read in using the bash source /etc/ipip.routes command (assuming you called the file with the routing commands /etc/ipip.routes) as illustrated. The source file must be in the NOS route command format.

• Note the use of the window argument on the route command. Setting this parameter to an appropriate value improves the performance of your radio link.

The new tunnel-munge script:

#!/bin/sh
#
# From: Ron Atkinson <n8fow@hamgate.cc.wayne.edu>
#
#  This script is basically the 'munge' script written by Bdale N3EUA
#  for the IPIP daemon and is modified by Ron Atkinson N8FOW. It's
#  purpose is to convert a KA9Q NOS format gateways route file
#  (usually called 'encap.txt') into a Linux routing table format
#  for the IP tunnel driver.
#
#        Usage: Gateway file on stdin, Linux route format file on stdout.
#               eg.  tunnel-munge < encap.txt > ampr-routes
#
# NOTE: Before you use this script be sure to check or change the
#       following items:
#
#     1) Change the 'Local routes' and 'Misc user routes' sections
#        to routes that apply to your own area (remove mine please!)
#     2) On the fgrep line be sure to change the IP address to YOUR
#        gateway Internet address. Failure to do so will cause serious
#        routing loops.
#     3) The default interface name is 'tunl0'. Make sure this is
#        correct for your system.
echo "#"
echo "# IP tunnel route table built by $LOGNAME on `date`"
echo "# by tunnel-munge script v960307."
echo "#"
echo "# Local routes"
echo "route add -net 44.xxx.xxx.xxx netmask 255.mmm.mmm.mmm dev sl0"
echo "#"
echo "# Misc user routes"
echo "#"
echo "# remote routes"
fgrep encap | grep "^route" | grep -v " XXX.XXX.XXX.XXX" | \
awk '{
        split($3, s, "/")
        split(s[1], n,".")
        if      (n[1] == "")    n[1]="0"
        if      (n[2] == "")    n[2]="0"
        if      (n[3] == "")    n[3]="0"

if      (n[4] == "")    n[4]="0"
if      (s[2] == "1")   mask="128.0.0.0"
else if (s[2] == "2")   mask="192.0.0.0"
else if (s[2] == "3")   mask="224.0.0.0"
else if (s[2] == "4")   mask="240.0.0.0"
else if (s[2] == "5")   mask="248.0.0.0"
else if (s[2] == "6")   mask="252.0.0.0"
else if (s[2] == "7")   mask="254.0.0.0"
else if (s[2] == "8")   mask="255.0.0.0"
else if (s[2] == "9")   mask="255.128.0.0"
else if (s[2] == "10")  mask="255.192.0.0"
else if (s[2] == "11")  mask="255.224.0.0"
else if (s[2] == "12")  mask="255.240.0.0"
else if (s[2] == "13")  mask="255.248.0.0"
else if (s[2] == "14")  mask="255.252.0.0"
else if (s[2] == "15")  mask="255.254.0.0"
else if (s[2] == "16")  mask="255.255.0.0"
else if (s[2] == "17")  mask="255.255.128.0"
else if (s[2] == "18")  mask="255.255.192.0"
else if (s[2] == "19")  mask="255.255.224.0"
else if (s[2] == "20")  mask="255.255.240.0"
else if (s[2] == "21")  mask="255.255.248.0"
else if (s[2] == "22")  mask="255.255.252.0"
else if (s[2] == "23")  mask="255.255.254.0"
else if (s[2] == "24")  mask="255.255.255.0"
else if (s[2] == "25")  mask="255.255.255.128"
else if (s[2] == "26")  mask="255.255.255.192"
else if (s[2] == "27")  mask="255.255.255.224"
else if (s[2] == "28")  mask="255.255.255.240"
else if (s[2] == "29")  mask="255.255.255.248"
else if (s[2] == "30")  mask="255.255.255.252"
else if (s[2] == "31")  mask="255.255.255.254"
else                    mask="255.255.255.255"
if (mask == "255.255.255.255")
printf "route add -host %s.%s.%s.%s gw %s dev tunl0\n"\
        ,n[1],n[2],n[3],n[4],$5
else
printf "route add -net %s.%s.%s.%s gw %s netmask %s dev tunl0\n"\
        ,n[1],n[2],n[3],n[4],$5,mask
}'
echo "#"
echo "# default the rest of amprnet via mirrorshades.ucsd.edu"
echo "route add -net 44.0.0.0 gw 128.54.16.18 netmask 255.0.0.0 dev tunl0"
echo "#"
echo "# the end"

21.3 AXIP encapsulated gateway configuration

Many Amateur Radio Internet gateways encapsulate AX.25, NetRom and Rose in addition to tcp/ip. Encapsulation of AX.25 frames within IP datagrams is described in RFC-1226 by Brian Kantor. Mike Westerhof wrote an implementation of an AX.25 encapsulation daemon for unix in 1991. The ax25-utils package includes a marginally enhanced version of it for Linux.

An AXIP encapsulation program accepts AX.25 frames at one end, looks at the destination AX.25 address to determine what IP address to send them to, encapsulates them in a tcp/ip datagram and then transmits them to the appropriate remote destination. It also accepts tcp/ip datagrams that contain AX.25 frames, unwraps them and processes them as if it had received them directly from an AX.25 port. To distinguish IP datagrams containing AX.25 frames from other IP datagrams which don’t, AXIP datagrams are coded with a protocol id of 4 (or 94 which is now deprecated). This process is described in RFC-1226.

The ax25ipd program included in the ax25-utils package presents itself as a program supporting a KISS interface across which you pass AX.25 frames, and an interface into the tcp/ip protocols. It is configured with a single configuration file called /etc/ax25/ax25ipd.conf.

21.3.1 AXIP configuration options.

The ax25ipd program has two major modes of operation. "digipeater" mode and "tnc" mode. In "tnc" mode the daemon is treated as though it were a kiss TNC, you pass KISS encapsulated frames to it and it will transmit them, this is the usual configuration. In "digipeater" mode, you treat the daemon as though it were an AX.25 digipeater. There are subtle differences between these modes.

In the configuration file you configure "routes" or mappings between destination AX.25 callsigns and the IP addresses of the hosts that you want to send the AX.25 packets too. Each route has options which will be explained later.

Other options that are configured here are

the tty that the ax25ipd daemon will open and its speed (usually one end of a pipe)

what callsign you want to use in "digipeater" mode

beacon interval and text

whether you want to encapsulate the AX.25 frames in IP datagrams or in UDP/IP datagrams. Nearly all AXIP gateways use IP encapsulation, but some gateways are behind firewalls that will not allow IP with the AXIP protocol id to pass and are forced to use UDP/IP. Whatever you choose must match what the tcp/ip host at the other end of the link is using.

21.3.2 A typical /etc/ax25/ax25ipd.conf file.
     #
     # ax25ipd configuration file for station floyd.vk5xxx.ampr.org
     #
     # Select axip transport. 'ip' is what you want for compatibility
     # with most other gateways.
     #
     socket ip

#
# Set ax25ipd mode of operation. (digi or tnc)
#
mode tnc
#
# If you selected digi, you must define a callsign.  If you selected
# tnc mode, the callsign is currently optional, but this may change
# in the future! (2 calls if using dual port kiss)
#
#mycall vk5xxx-4
#mycall2 vk5xxx-5
#
# In digi mode, you may use an alias. (2 for dual port)
#
#myalias svwdns
#myalias2 svwdn2
#
# Send an ident every 540 seconds ...
#
#beacon after 540
#btext ax25ip -- tncmode rob/vk5xxx -- Experimental AXIP gateway
#
# Serial port, or pipe connected to a kissattach in my case
#
device /dev/ttyq0
#
# Set the device speed
#
speed 9600
#
# loglevel 0 - no output
# loglevel 1 - config info only
# loglevel 2 - major events and errors
# loglevel 3 - major events, errors, and AX25 frame trace
# loglevel 4 - all events
# log 0 for the moment, syslog not working yet ...
#
loglevel 2
#
# If we are in digi mode, we might have a real tnc here, so use param to
# set the tnc parameters ...
#
#param 1 20
#
# Broadcast Address definition. Any of the addresses listed will be forwarded
# to any of the routes flagged as broadcast capable routes.
#
broadcast QST-0 NODES-0
#
# ax.25 route definition, define as many as you need.
# format is route (call/wildcard) (ip host at destination)

     # ssid of 0 routes all ssid's
     #
     # route <destcall> <destaddr> [flags]
     #
     # Valid flags are:
     #         b  - allow broadcasts to be transmitted via this route
     #         d  - this route is the default route
     #
     route vk2sut-0 44.136.8.68 b
     route vk5xxx 44.136.188.221 b
     route vk2abc 44.1.1.1
     #
     #

21.3.3 Running ax25ipd

Create your /etc/ax25/axports entry:
          # /etc/ax25/axports
          #
          axip    VK2KTJ-13       9600    256     AXIP port
          #

Run the kissattach command to create that port:

          /usr/sbin/kissattach /dev/ptyq0 axip

Run the ax25ipd program:

          /usr/sbin/ax25ipd &

Test the AXIP link:

          call axip vk5xxx

21.3.4 Some notes about the routes and route flags

The "route" command is where you specify where you want your AX.25 packets encapsulated and sent to. When the ax25ipd daemon receives a packet from its interface, it compares the destination callsign with each of the callsigns in its routing table. If if finds a match then the ax.25 packet is encapsulated in an IP datagram and then transmitted to the host at the specified IP address.

There are two flags you can add to any of the route commands in the ax25ipd.conf file. The two flags are:

b

traffic with a destination address matching any of those on the list defined by the "broadcast" keyword should be transmitted via this route.

d

any packets not matching any route should be transmitted via this route.

The broadcast flag is very useful, as it enables informations that is normally destined for all stations to a number of AXIP destinations. Normally axip routes are point-to-point and unable to handle ’broadcast’ packets.

21.4 Linking NOS and Linux using a pipe device

Many people like to run some version of NOS under Linux because it has all of the features and facilities they are used to. Most of those people would also like to have the NOS running on their machine capable of talking to the Linux kernel so that they can offer some of the linux capabilities to radio users via NOS.

Brandon S. Allbery, KF8NH, contributed the following information to explain how to interconnect the NOS running on a Linux machine with the kernel code using the Linux pipe device.

Since both Linux and NOS support the slip protocol it is possible to link the two together by creating a slip link. You could do this by using two serial ports with a loopback cable between them, but this would be slow and costly. Linux provides a feature that many other Unix-like operating systems provide called ‘pipes’. These are special pseudo devices that look like a standard tty device to software but in fact loopback to another pipe device. To use these pipes the first program must open the master end of the pipe, and the open then the second program can open the slave end of the pipe. When both ends are open the programs can communicate with each other simply by writing characters to the pipes in the way they would if they were terminal devices.

To use this feature to connect the Linux Kernel and a copy of NOS, or some other program you first must choose a pipe device to use. You can find one by looking in your /dev directory. The master end of the pipes are named: ptyq[1-f] and the slave end of the pipes are known as: ttyq[1-f]. Remember they come in pairs, so if you select /dev/ptyqf as your master end then you must use /dev/ttyqf as the slave end.

Once you have chosen a pipe device pair to use you should allocate the master end to you linux kernel and the slave end to the NOS program, as the Linux kernel starts first and the master end of the pipe must be opened first. You must also remember that your Linux kernel must have a different IP address to your NOS, so you will need to allocate a unique address for it if you haven’t already.

You configure the pipe just as if it were a serial device, so to create the slip link from your linux kernel you can use commands similar to the following:

     # /sbin/slattach -s 38400 -p slip /dev/ptyqf &
     # /sbin/ifconfig sl0 broadcast 44.255.255.255 pointopoint 44.70.248.67 /
             mtu 1536 44.70.4.88
     # /sbin/route add 44.70.248.67 sl0
     # /sbin/route add -net 44.0.0.0 netmask 255.0.0.0 gw 44.70.248.67

In this example the Linux kernel has been given IP address 44.70.4.88 and the NOS program is using IP address 44.70.248.67. The route command in the last line simply tells your linux kernel to route all datagrams for the amprnet via the slip link created by the slattach command. Normally you would put these commands into your /etc/rc.d/rc.inet2 file after all your other network configuration is complete so that the slip link is created automatically when you reboot. Note: there is no advantage in using cslip instead of slip as it actually reduces performance because the link is only a virtual one and occurs fast enough that having to compress the headers first takes longer than transmitting the uncompressed datagram.

To configure the NOS end of the link you could try the following:

     # you can call the interface anything you want; I use "linux" for convenience.
     attach asy ttyqf - slip linux 1024 1024 38400
     route addprivate 44.70.4.88 linux

These commands will create a slip port named ‘linux’ via the slave end of the pipe device pair to your linux kernel, and a route to it to make it work. When you have started NOS you should be able to ping and telnet to your NOS from your Linux machine and vice versa. If not, double check that you have made no mistakes especially that you have the addresses configured properly and have the pipe devices around the right way.

22 Where do I find more information about .... ?

Since this document assumes you already have some experience with packet radio and that this might not be the case I’ve collected a set of references to other information that you might find useful.

22.1 Packet Radio

You can get general information about Packet Radio from these sites:

American Radio Relay League <http://www.arrl.org/>,

Radio Amateur Teleprinter Society <http://www.rats.org/>

Tucson Amateur Packet Radio Group <http://www.tapr.org/>

22.2 Protocol Documentation

AX.25, NetRom - Jonathon Naylor has collated a variety of documents that relate to the packet radio protocols themselves. This documentation has been packaged up into ax25-doc-1.0.tar.gz <ftp:

     //ftp.pspt.fi/pub/ham/linux/ax25/ax25-doc-1.0.tar.gz>

22.3 Hardware Documentation

Information on the PI2 Card is provided by the Ottawa Packet Radio Group <http://hydra.

     carleton.ca/>.

Information on Baycom hardware is available at the Baycom Web Page <http://www.baycom.de/>.

23 Discussion relating to Amateur Radio and Linux.

There are various places that discussion relating to Amateur Radio and Linux take place. They take place in the comp.os.linux.* newsgroups, they also take place on the HAMS list on vger.rutgers.edu. Other places where they are held include the tcp-group mailing list at ucsd.edu (the home of amateur radio TCP/IP discussions), and you might also try the #linpeople channel on the linuxnet irc network.

To join the Linux linux-hams channel on the mail list server, send mail to:

     Majordomo@vger.rutgers.edu

with the line:

     subscribe linux-hams

in the message body. The subject line is ignored.

The linux-hams mailing list is archived at:

zone.pspt.fi <http://zone.pspt.fi/archive/linux-hams/> and zone.oh7rba.ampr.org <http://zone.

oh7rba.ampr.org/archive/linux-hams/>. Please use the archives when you are first starting, because many common questions are answered there.

To join the tcp-group send mail to:

     listserver@ucsd.edu

with the line:

     subscribe tcp-group

in the body of the text.

Note: Please remember that the tcp-group is primarily for discussion of the use of advanced protocols, of which TCP/IP is one, in Amateur Radio. Linux specific questions should not ordinarily go there.

24 Acknowledgements.

The following people have contributed to this document in one way or another, knowingly or unknowingly. In no particular order (as I find them): Jonathon Naylor, Thomas Sailer, Joerg Reuter, Ron Atkinson, Alan Cox, Craig Small, John Tanner, Brandon Allbery, Hans Alblas, Klaus Kudielka, Carl Makin.

25 Copyright.

The AX25-HOWTO, information on how to install and configure some of the more important packages providing AX25 support for Linux. Copyright (c) 1996 Terry Dawson.

This program is free software; you can redistribute it and/or modify it under the terms of the GNU General Public License as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version.

This program is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE. See the GNU General Public License for more details.

You should have received a copy of the GNU General Public License along with this program; if not, write to the:

Free Software Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.